### v1

# Should we pre-train a decoder in contrastive learning for dense prediction tasks?

Abstract
Contrastive learning in self-supervised settings primarily focuses on pre-training encoders, while decoders are typically introduced and trained separately for downstream dense prediction tasks. This conventional approach, however, overlooks the potential benefits of jointly pre-training both the encoder and decoder. In this paper, we propose DeCon: a framework-agnostic adaptation to convert an encoder-only self-supervised learning (SSL) contrastive approach to an efficient encoder-decoder framework that can be pre-trained in a contrastive manner. We first update the existing architecture to accommodate a decoder and its respective contrastive loss. We then introduce a weighted encoder-decoder contrastive loss with non-competing objectives that facilitates the joint encoder-decoder architecture pre-training. We adapt two established contrastive SSL frameworks tailored for dense prediction tasks, achieve new state-of-the-art results in COCO object detection and instance segmentation, and match state-of-the-art performance on Pascal VOC semantic segmentation. We show that our approach allows for pre-training a decoder and enhances the representation power of the encoder and its performance in dense prediction tasks. This benefit holds across heterogeneous decoder architectures between pre-training and fine-tuning and persists in out-of-domain, limited-data scenarios.
**footnotetext: Equal contributions for this work. Co-first authors are ordered in alphabetical order.
1 Introduction
The increasing demand for efficient Deep Learning (DL) methods arises from their ability to address complex tasks. While traditional approaches like supervised learning can yield powerful, high-performing models, they depend heavily on large volumes of high-quality labeled data. However, obtaining such annotations is often tedious, expensive, and sometimes impractical, even when abundant data is available. To overcome this bottleneck, one can benefit from a pre-trained model. Pre-trained models are usually models trained from scratch on a large scale dataset and allow for better performance on downstream tasks with smaller datasets. Typically, an encoder is pre-trained to capture a representation of the input data, which can then be fine-tuned for downstream tasks such as classification or segmentation. During the downstream fine-tuning phase, randomly initialized layers are typically added on top of the encoder, and the complete architecture is retrained using a supervised learning approach on part of or on a full annotated dataset. ImageNet [15] is a large and widely utilized dataset with annotated classes. However, models pre-trained on ImageNet classification are not the best fit for every downstream tasks, especially when the downstream objective is not classification. Since creating large-scale annotated dataset for every tasks presents numerous challenges, Self-Supervised Learning (SSL) has emerged as a promising alternative, enabling models to learn from the vast amount of unlabeled data and perform effectively on various downstream tasks. These pre-trained models can reach downstream performance close to a fully supervised approach using much smaller to no labeled datasets. Notable approaches such as SimCLR [7], VicReg [2], and MoCo [22] showed promising results by leveraging contrastive learning for downstream classification. However, most SSL frameworks primarily target classification downstream tasks. Some efforts such as PixCon [37] or DenseCL [45], focus on dense prediction tasks like segmentation or object detection. In all of these efforts, authors only train the encoders of the DL architectures and focus on incorporating a local loss with existing classification-targeted SSL approaches to adapt with dense tasks.
In this work, we propose a Decoder-aware contrastive learning (DeCon) approach, a framework agnostic encoder-decoder adaptation to pre-train an encoder jointly with a decoder, improving its representation power and efficiently preparing it to be combined with a decoder in downstream tasks. Our contributions are as follows.
  • We propose DeCon-SL, a single-level joint encoder-decoder adaptation of contrastive dense SSL frameworks. We show that jointly pre-training the encoder with a decoder enhances the representation power of the encoder and improves downstream dense prediction performance, even without transferring the pre-trained decoder.
  • We improve on this adaptation by introducing DeCon-ML, which extends DeCon-SL with a multi-level decoder loss function and an encoder level-wise channel dropout to promote a comprehensive feature usage of the encoder. It achieves state-of-the-art results on COCO object detection and instance segmentation and matches top performance on Pascal VOC semantic segmentation.
  • We demonstrate that DeCon improves the encoder’s downstream performance compared to encoder-only frameworks, even when pre-training and fine-tuning decoders differ.
  • We show that DeCon’s efficiency persists in out-of-domain downstream tasks and provides even greater benefits in scenarios with limited annotated fine-tuning data.
2 Related works
Self-Supervised pre-training: In SSL, deep learning models are trained on pretext tasks where the data itself provides supervision, which allow models to learn meaningful representations without any manual annotations. Early SSL works focused on designing these tasks, such as predicting missing parts of an image (image inpainting) [38], context prediction [16], solving jigsaw puzzles [34], colorizing grayscale images [52], and predicting rotations [20]. These methods aimed to create surrogate objectives that encourage models to develop feature representations, which could later be transferred to downstream tasks. However, these methods often produced task-specific representations with limited generalizability and restricted scalability.
Recent methods prioritize generalizability and move away from hand-crafted pretext tasks. These approaches can be broadly categorized into generative and contrastive methods. Generative approaches in SSL aim to learn meaningful representations by modeling the underlying data distribution. This is typically achieved through reconstruction-based methods, where the model predicts corrupted parts of the input [23, 1, 17, 51], or adversarial methods, where a generative model learns to synthesize realistic samples while an encoder captures high-level features [48, 26]. Contrastive methods, which are the focus of this research, learn data representations by maximizing the similarity between features extracted from two augmented views of the same image. To avoid model collapse—a phenomenon where the model maps all inputs to the same representation—negative pairs [22, 8], stop-gradient mechanisms [9], variance regularization [2, 3], momentum contrast [22] or other architectural and optimization techniques are employed [5].
Dense pre-training: Even though most of the aforementioned frameworks can be used for downstream segmentation or detection, these models are not designed specifically for pixel-level, i.e., dense downstream tasks. Consequently, developing SSL approaches that generalize well to dense downstream tasks, such as the segmentation task, is an active area of research [46, 45, 44, 3, 35]. The main idea behind these approaches is to shift from a global image similarity to a local similarity. Local similarity can be enforced based on the location of pixels in an image [50, 44, 35], local pixel features [45], or both [29, 3, 37]. Local similarity can also be enforced in a region-based manner [19, 24, 49, 46]. These methods outperform those based solely on image-level global similarity in downstream dense tasks.
Encoder-Decoder pre-training: Encoder-decoder architectures with skip connections between the encoder and decoder layers (e.g., U-Net-like models) have achieved state-of-the-art results in many segmentation tasks [40, 27, 53]. A common approach for pre-training such models in a self-supervised manner is image reconstruction [30, 4, 6]. However, the presence of skip connections allows a substantial amount of information to bypass the encoder’s final layer—often referred to as the bottleneck layer–resulting in lower-quality embeddings generated by the encoder. Our proposed approach preserves skip connections while addressing the issue of inferior representations generated by the decoder. Encoder-only contrastive learning frameworks are built solely around an encoder, incorporating components such as predictors and projectors. However, they do not include a decoder, even though a randomly initialized decoder is later attached to the pre-trained encoder in downstream dense tasks to form a complete architecture. While some generative SSL frameworks use a decoder for pre-training [14, 38, 23], to the best of our knowledge, no work has explored pre-training vision models with both encoder and decoder loss in a contrastive learning framework to learn a data representation.
Figure 1: A: DeCon-SL. Instead of a classical encoder-only pre-training, a decoder is pre-trained alongside the encoder. We mirror the encoder loss at the decoder level and optimize the architecture using a weighted sum of the encoder and decoder losses. B: DeCon-ML. Instead of computing the decoder loss at a single level, it is calculated across multiple levels (four in this figure). Additionally, a channel-wise dropout is applied at the output of each encoder level before it is passed through the skip connection to the decoder.
3 Methods
We introduce two architectural adaptations, as depicted in  Fig. 1: (A) a single-level decoder loss (DeCon-SL) and (B) a multi-level decoder loss (DeCon-ML).
3.1 DeCon-SL
Given a vision SSL framework that pre-trains an encoder using a teacher-student architecture [25], we retain the encoders and any existing “auxiliary layers” (e.g., predictors, projection heads), and add decoders and their corresponding auxiliary layers for both teacher and student networks. For all models, the auxiliary layers are designed to match those of the encoder, with adjustments applied to align with the decoder’s output feature size. We obtain two losses, one is computed from the encoder features while the other one is computed from the decoder features. Consequently, the loss function is defined as follows:
$Loss=\alpha\times L_{enc}+(1-\alpha)\times L_{dec}$ (1)
where, $\alpha$ represents the contribution of the encoder loss in the objective function.
The DeCon-SL architecture is illustrated in Fig. 1 A, where $f_{\theta}$ and $f_{\epsilon}$ represent teacher and student encoders, respectively. $aux^{enc}_{\theta}$ and $aux^{enc}_{\epsilon}$ are the SSL framework’s auxiliary layers of the teacher and student encoders, respectively. Auxiliary layers can include any projector, predictor heads, or any layers used in the loss definition of the original SSL framework. $g_{\theta}$ and $g_{\epsilon}$ are the decoders of the teacher and student networks, and $aux^{dec}_{\theta}$ and $aux^{dec}_{\epsilon}$ are the auxiliary layers feeding from their respective decoder. $g_{\theta}$ and $g_{\epsilon}$ take the output of encoder $f_{\theta}$ and $f_{\epsilon}$ as input, respectively. Auxiliary layers $aux^{enc}_{\theta}$ and $aux^{enc}_{\epsilon}$ are responsible for generating a representation from the encoder output, that is used to compute a contrastive loss for the encoders $L^{encoder}$. The contrastive loss, $L^{decoder}$ is calculated from the representation obtained from the auxiliary layers of the decoders, i.e., $aux^{dec}_{\theta}$ and $aux^{dec}_{\epsilon}$. $f_{\epsilon}$ and $g_{\epsilon}$ are updated using the exponential moving average (EMA) of $f_{\theta}$ and $g_{\theta}$, respectively.
3.2 DeCon-ML
To address the challenge posed by skip connections between the encoder and decoder layers (as mentioned in the Related works section), we introduce a multi-layer contrastive decoder loss, consisting of a single loss for the encoder and multiple loss components for the decoder along with channel dropout and deep supervision, which we describe in detail below. Figure 1 B illustrates the DeCon-ML adaptation of a given SSL framework.
Channel dropout: The dropout layer is typically used to prevent overfitting, improve generalization, and act as an implicit ensemble [41, 43]. However, we extend its use by applying dropout to channels transferred through skip connections between the encoder and decoder. This prevents the model from relying heavily on specific features shared through skip connections and enforces full encoder training, leading to richer data representations at the bottleneck layer of the encoder. In DeCon-ML, we apply channel dropout to the output layers of the encoder at the different levels when passing these outputs to the decoder. No channel is zeroed-out when the inputs pass through the encoder. This channel dropout promotes a more comprehensive utilization of the encoder parameters and enhances representation power at various encoder levels while preserving the input information content as it passes through the encoder to its bottleneck.
Decoder deep supervision: An encoder-decoder SSL framework allows us to retrieve meaningful features from multiple levels of the decoder. We can then pre-train the architecture with multiple losses using decoder deep supervision. If the decoder is connected to the encoder at multiple levels through lateral/skip connections, decoder deep supervision strengthens the representation power of the encoder at different levels [39]. Since a decoder passes the encoder bottleneck representation in a bottom-up manner to its upper level layers, the final encoder bottleneck representation remains a non-negligible objective of the training since it impacts all decoder level losses. In contrast, applying deep supervision at multiple encoder stages in an encoder-only SSL framework might prevent the creation of a strong semantic representation at the encoder bottleneck. To perform the decoder deep supervision, we add auxiliary layers to each decoder level, and a loss is computed at each level between the teacher and student representations. When performing decoder deep supervision, the encoder loss will remain the same as Eq. 1, but the decoder loss will be updated as follows:
$Loss=$ $\alpha\times L_{enc}+(1-\alpha)\times L_{dds}$ (2)
$L_{dds}=$ $\frac{1}{j}\times\sum_{i=1}^{j}L_{dec_{i}}$ (3)
where $j$ is the number of decoder levels at which a loss will be calculated and $L_{dec_{i}}$ is the loss computed at level $i$. The final decoder loss in Eq. 2 is thus an average of the losses at the different levels.
3.3 Implementation details
Architecture: Following common practice in contrastive SSL approaches literature, we use a ResNet-50 encoder [21] as a backbone, a Feature Pyramid Network (FPN) [32] for downstream detection and instance segmentation, and a Fully Convolutional Network (FCN) [33] for downstream semantic segmentation [45, 22, 46] tasks. We implement FCN and skip-connected FPN as two decoder architectures in the DeCon framework adaptation. We follow the mmseg [11] FCN implementation used in SlotCon [46] and customize FPN implementation from Detectron2 [47] version 0.6. We use DeCon to adapt the SlotCon [46] and DenseCL [45] frameworks. More implementation details are available in Supplementary Material Section 1.
For SlotCon, the student and teacher encoders, and every auxiliary layers—i.e., the respective projectors and semantic groupings, the slot predictor and the prototypes—are kept unchanged. Two decoders, a student and a teacher decoder (either FCN or FPN), that take the output of their corresponding encoder as input are added to the framework. The auxiliary layers are replicated at the decoder level. In the case of decoder deep supervision with FPN decoders, auxiliary layers are replicated independently at the four decoder levels. No deep supervision is performed in the DeCon adaptation of SlotCon with an FCN decoder. The decoder projectors differ from the encoder projectors only in input channels: 2048 for the encoder and 256 for the FPN and FCN decoders. We follow SlotCon by updating the teacher branch and the encoder and decoder slot predictors with Backpropagation, the student branch with an EMA of the first one, and the prototypes centers with an EMA of the previous batch centers seen during training. Our implementation is based on the official SlotCon repository. Detailed figures of SlotCon DeCon-SL and DeCon-ML adaptations are available in the Supplementary Material Section 1.
For DenseCL, the two encoders, along with their respective auxiliary layers—i.e., the global and dense projection heads—are kept the same as well as the independent dictionaries for global and dense losses. Additionally, we introduce two decoders, which take their respective encoders’ outputs as inputs. We also replicate the encoder’s dense and global projection heads and dictionaries separately for the two decoders. As for SlotCon, the input dimension of the decoder projectors is 256 instead of 2048 for the encoder. The encoders and decoders dictionaries size is set to 16384. We follow DenseCL by updating the first branch of the architecture through Backpropagation, while the second branch is updated using an EMA of the first one. Due to extensive memory requirements of replicating dictionaries at each decoder level, we do not conduct decoder deep supervision or channel dropout (DeCon-ML) experiments with DenseCL. We adapt DenseCL implementation from mmselfsup package [12] version 1.0.0.
Pre-training setup: We pre-train DeCon on COCO 2017 dataset [31]. The augmentations and all the hyperparameters—learning rate, weight decay, schedulers, batch size, input shape, number of epochs, momentum, and so on—are kept the same as used in the original SSL framework: SlotCon or DenseCL. The SlotCon models are pre-trained for 800 epochs using a LARS optimizer, with a total batch size of 512 across all available GPUs. The base learning rate is 1.0 scaled with the batch size and updated with a cosine learning rate decay schedule using weight decay of $10^{-5}$ and a warm-up period of 5 epochs. The pre-trainings using FPN decoders are done using eight NVIDIA H100 GPUs (80GB each). The pre-trainings without decoder and with FCN decoders are conducted on one 80GB NVIDIA A100 GPU. The DenseCL models are pre-trained for 800 epochs using an SGD optimizer with a total batch size of 256. The base learning rate is 0.3 and is updated using a cosine learning rate decay schedule (0.0001 weight decay, 0.9 momentum). The pre-training experiments are conducted with eight NVIDIA A100 GPUs (40GB each).
To assess the DeCon framework on non scene-centric data, we pre-train SlotCon and a DeCon-SL adaptation of SlotCon with an FCN decoder on two small-scale datasets: ISIC 2017 and REFUGE described in Sec. 3.4. For ISIC dataset, we pre-train for 800 epochs with a batch size of 256 and a base learning rate of 1.0 linearly scaled with batch size. For REFUGE, we pre-train for 2400 epochs with a batch size of 192 and the same learning rate as for ISIC. These pre-trainings are performed with one NVIDIA A6000 GPU with 48GB of memory.
3.4 Evaluation protocol
We evaluate the different pre-trained models on object detection, semantic segmentation, and instance segmentation as downstream tasks. We either transfer the pre-trained encoder or both the encoder and decoder for each fine-tuning experiment. All fine-tuning results reported in this paper are averaged over three independent runs. The pre-trainings, however, are only run once. All fine-tuning experiments related to the same dataset are run on a similar platform with the same hardware. We perform object detection and instance segmentation using Detectron2 package [47] version 0.6, whereas mmsegmentation [11] version 0.30.0 is used for semantic segmentation tasks.
Object detection and instance segmentation: We perform object detection on Pascal VOC dataset [18] and object detection and instance segmentation on COCO dataset [31]. We fine-tune a Faster R-CNN with a FPN C4 backbone using VOC trainval07+12 dataset for 24000 iterations and evaluate on the VOC test2007 set following previous works [22, 50]. For COCO object detection and instance segmentation, we fine-tune a Mask R-CNN with FPN backbone on the COCO train2017 dataset with 90000 iterations and evaluate on COCO val2017 split following SlotCon [46]. We initialize the network either with a pre-trained ResNet-50 encoder or with both a ResNet-50 encoder and a pre-trained FPN decoder, and fine-tune end-to-end in all cases. All the aforementioned experiments are performed using one NVIDIA A100 GPU with 80GB memory.
Semantic segmentation on in-domain datasets: We perform semantic segmentation on Pascal VOC [18] and Cityscapes [13] datasets that contain scene-centric data following SlotCon [46] and PixCon [37]. A ResNet-50 encoder with a two-layer FCN decoder, similar to the FCN decoder used during pre-training, is used. For Pascal VOC, we train on VOC train_aug2012 set for 30000 iterations with a batch size of 16 and evaluate on VOC val2012 set. For Cityscapes, we fine tune on the train_fine set for 90000 iterations with a batch size of 16 and evaluate on the val_fine set. We initialize the network either with a pre-trained ResNet-50 encoder or with both the ResNet-50 encoder and the FCN decoder pre-trained and fine-tune end-to-end in all cases. All the aforementioned experiments are performed using one NVIDIA A100 GPU with 80GB (Cityscapes) or 40GB (Pascal VOC) memory.
Semantic segmentation on out-of-domain datasets: We also evaluate generalizability to out-of-domain datasets for downstream tasks on different dataset sizes, considering 5, 25, and 100% randomly selected training samples. We fine-tune a ResNet-50 encoder coupled with an FCN decoder for semantic segmentation on REFUGE [36] and ISIC 2017 [10] dataset. REFUGE dataset contains 1200 retinal images in total representing 3 different classes: optic disk, optic cup, and background, divided into 400 training, 400 validation, and 400 testing samples. We train the ResNet-50-FCN network end-to-end for 80000 iterations with a batch size of 16 on the training set using one NVIDIA A100 GPU (40GB). The learning rate starts at 0.01 and decreases with a polynomial scheduler with a power of 0.9. We evaluate the model on the test set using the last iteration checkpoint and report the results.
ISIC 2017 is a skin-lesion segmentation dataset which consists of 2000 training, 150 validation, and 600 testing images. We fine-tune the same ResNet-50-FCN architecture on the training set for 24000 iterations on one NVIDIA V100 with 32GB of memory. Hyper parameters are the same as for REFUGE. We evaluate on the validation set every 500 iterations, select the best checkpoint and report the results of this checkpoint on the test set.
Table 1: Performance of DeCon-SL adaptation ($\alpha=0.5$) of SlotCon and DenseCL using FPN and FCN as options for decoders as well as baseline approaches. COCO dataset was used for pre-training. The same decoder architecture was used for pre-training and fine-tuning. The architecture was then fine-tuned on various segmentation and object detection tasks. All fine-tuning results were averaged over three runs. Up and down arrows indicate better and worst performance compared to baseline, respectively.
Framework Adaptation Decoder Loss Transfer Object Detection Object Detection Instance Seg. Semantic Seg.
VOC COCO COCO VOC City
$L_{enc}$ $L_{dec}$ enc dec AP AP50 AP75 AP AP50 AP75 AP AP50 AP75 mIoU mIoU
SlotCon None None 55.16 81.76 61.23 40.81 60.95 44.37 36.80 57.98 39.54 71.50 75.95
DeCon-SL FPN 55.57  $\uparrow$ 81.75  $\downarrow$ 61.76  $\uparrow$ 40.94  $\uparrow$ 60.97  $\uparrow$ 44.82  $\uparrow$ 36.92  $\uparrow$ 58.03  $\uparrow$ 39.80  $\uparrow$ - -
55.40  $\uparrow$ 81.48  $\downarrow$ 61.21  $\downarrow$ 41.05  $\uparrow$ 61.34  $\uparrow$ 44.85  $\uparrow$ 37.03  $\uparrow$ 58.33  $\uparrow$ 39.86  $\uparrow$ - -
FCN - - - - - - - - - 72.42  $\uparrow$ 75.67  $\downarrow$
- - - - - - - - - 72.58  $\uparrow$ 76.00  $\uparrow$
DenseCL None None 54.50 80.47 59.97 39.60 59.30 43.29 35.65 56.49 38.25 71.11 75.87
DeCon-SL FPN 54.71  $\uparrow$ 80.65  $\uparrow$ 60.44  $\uparrow$ 39.60 59.36  $\uparrow$ 43.31  $\uparrow$ 35.74  $\uparrow$ 56.45  $\downarrow$ 38.40  $\uparrow$ - -
54.41  $\downarrow$ 80.24  $\downarrow$ 60.01  $\uparrow$ 39.55  $\downarrow$ 59.41  $\uparrow$ 43.15  $\downarrow$ 35.71  $\uparrow$ 56.44  $\downarrow$ 38.34  $\uparrow$ - -
FCN - - - - - - - - - 71.31  $\uparrow$ 76.04  $\uparrow$
- - - - - - - - - 71.59  $\uparrow$ 76.17  $\uparrow$
4 Results
DeCon-SL: Table 1 shows the results of DeCon-SL adaptation. DeCon pre-trained components outperformed baseline models in downstream object detection and segmentation when only the encoder was transferred, except for Cityscapes semantic segmentation with SlotCon. The transfer of both a pre-trained encoder and an FCN decoder always improved the performance further compared to transferring the pre-trained encoder only. Transferring the encoder and the FPN decoder pre-trained with DeCon-SL adaptation of DenseCL did not always result in improving model performance for downstream tasks.
DeCon-ML: Table 2 shows the results of DeCon-ML adaptation. We can see that adding the decoder deep supervision and a channel dropout probability of 0.5 during pre-training always enhanced fine-tuning performance compared to baseline approach (SlotCon) and DeCon-SL adaptation in all object detection and instance segmentation tasks.
Table 2: Performance of DeCon-ML adaptation of SlotCon with $\alpha=0.5$ and $dropout=0.5$ pre-trained on COCO dataset. We used an FPN as a DeCon decoder during pre-training and fine-tuning. We averaged all fine-tuning results over three runs. Arrows indicate performance gain or loss compared to baseline.
Frame work Adaptation Loss Transfer Object Detection Object Detection Instance Segmentation
VOC COCO COCO
$L_{enc}$ $L_{dec}$ $L_{dds}$ enc dec AP AP50 AP75 AP AP50 AP75 AP AP50 AP75
SlotCon None 55.16 81.76 61.23 40.81 60.95 44.37 36.80 57.98 39.54
DeCon-SL 55.57  $\uparrow$ 81.75  $\downarrow$ 61.76  $\uparrow$ 40.94  $\uparrow$ 60.97  $\uparrow$ 44.82  $\uparrow$ 36.92  $\uparrow$ 58.03  $\uparrow$ 39.80  $\uparrow$
55.40  $\uparrow$ 81.48  $\downarrow$ 61.21  $\downarrow$ 41.05  $\uparrow$ 61.34  $\uparrow$ 44.85  $\uparrow$ 37.03  $\uparrow$ 58.33  $\uparrow$ 39.86  $\uparrow$
DeCon-ML 55.60  $\uparrow$ 81.84  $\uparrow$ 61.83  $\uparrow$ 41.10  $\uparrow$ 61.37  $\uparrow$ 45.01  $\uparrow$ 37.05  $\uparrow$ 58.46  $\uparrow$ 39.87  $\uparrow$
55.57  $\uparrow$ 81.77  $\uparrow$ 61.76  $\uparrow$ 41.15  $\uparrow$ 61.38  $\uparrow$ 44.89  $\uparrow$ 37.05  $\uparrow$ 58.43  $\uparrow$ 39.88  $\uparrow$
Impact of the channel dropout: Impact of channel dropout is highlighted in Tab. 3. The decoder deep supervision on its own did not always improve downstream performance when only transferring the encoder in DeCon-ML. When transferring both encoder and decoder, downstream performance improved on Pascal VOC but not on COCO, which was the dataset used for pre-training. When the decoder deep supervision was coupled with channel dropout, performance improved compared to DeCon-SL. The channel dropout probability impacted the downstream performance differently when transferring only the encoder or both encoder and decoder. A channel dropout probability of 0.5 led to the best performance when transferring the encoder only. When transferring both encoder and decoder, a lower channel dropout probability of 0.25 led to the best performance on Pascal VOC object detection, and COCO instance segmentation. Comparing the performance of DeCon-ML adaptation with an encoder-only pre-training, we observed improvements in all scenarios regardless of dropout values.
Table 3: Impact of channel dropout probability on DeCon-ML. We pre-trained a DeCon-ML adaptation of SlotCon framework on COCO dataset with different channel dropout probabilities. We used an FPN as a DeCon decoder during pre-training and fine-tuning. We averaged all fine-tuning results over three runs.
Transfer SSL Loss $\alpha$ Dropout VOC Object Detection COCO Object Detection COCO Instance Seg.
$L_{enc}$ $L_{dec}$ $L_{dds}$ AP AP50 AP75 AP AP50 AP75 AP AP50 AP75
Enc SlotCon 1 0 55.16 81.76 61.23 40.81 60.95 44.37 36.80 57.98 39.54
DeCon-SL 0.5 0 55.57 81.75 61.76 40.94 60.97 44.82 36.92 58.03 39.80
0.5 55.20 81.56 61.33 40.90 61.23 44.64 36.91 58.30 39.67
DeCon-ML 0 55.52 81.67 62.01 40.85 60.95 44.67 36.83 58.12 39.55
0.25 55.54 81.88 61.10 40.95 61.13 44.88 36.97 58.22 39.76
0.5 55.60 81.84 61.83 41.10 61.37 45.01 37.05 58.46 39.87
0.75 55.29 81.64 61.36 41.02 61.27 44.96 37.05 58.33 39.85
Enc + Dec DeCon-SL 0.5 0 55.40 81.48 61.21 41.05 61.34 44.85 37.03 58.33 39.86
DeCon-ML 0 55.70 81.89 61.69 40.97 61.37 44.66 37.08 58.45 39.88
0.25 55.79 81.79 61.93 41.06 61.36 44.99 37.13 58.52 39.86
0.5 55.57 81.77 61.76 41.15 61.38 44.89 37.05 58.43 39.88
0.75 55.37 81.57 61.48 41.11 61.45 44.91 37.13 58.55 39.99
Impact of encoder loss weight in DeCon: As shown in Tab. 4, the encoder loss weight $\alpha$ played a crucial role in downstream performance. Decreasing $\alpha$ down to 0 for DeCon-ML pre-training and down to 0.25 for DeCon-SL pre-training increased the fine-tuning performance. This pattern was also observed when transferring both the encoder and the decoder. The downstream performance of DeCon adaptation was higher than that of encoder-only pre-training (baseline, $\alpha=1$) for object detection and instance segmentation tasks, whether transferring only the encoder or both the encoder and decoder. For Cityscapes semantic segmentation, no pattern was observed regarding the relationship between the value of $\alpha$ and model performance.
Table 4: Impact of the encoder weight $\alpha$. We pre-trained a DeCon-ML adaptation of SlotCon with an FPN decoder ($dropout=0.5$) and a DeCon-SL adaptation of SlotCon with an FCN decoder on COCO dataset. We averaged all fine-tuning results over three runs.
Transfer $\alpha$ SSL strategy Object Detection Object Detection Instance Seg. SSL strategy Semantic Seg.
VOC COCO COCO VOC City
AP AP50 AP75 AP AP50 AP75 AP AP50 AP75 mIoU mIoU
Enc 1 SlotCon 55.16 81.76 61.23 40.81 60.95 44.37 36.80 57.98 39.54 SlotCon 71.50 75.95
0.75 DeCon-ML-FPN 0.5 Dropout ($L_{enc}+L_{dds}$) 55.33 81.68 61.34 40.98 61.28 44.76 36.95 58.34 39.72 DeCon-SL-FCN ($L_{enc}+L_{ds})$ 71.98 76.18
0.5 55.60 81.84 61.83 41.10 61.37 45.01 37.05 58.46 39.87 72.42 75.62
0.25 55.69 81.98 62.05 41.05 61.19 44.85 37.08 58.29 39.70 73.01 76.21
0 55.70 81.91 62.07 41.18 61.38 44.91 37.12 58.35 39.94 72.17 75.71
Enc + Dec 0.75 55.46 81.51 61.35 40.96 61.10 44.87 36.90 58.07 39.73 72.80 76.06
0.5 55.57 81.77 61.76 41.15 61.38 44.89 37.05 58.43 39.88 72.58 75.97
0.25 55.67 81.91 61.62 41.17 61.41 44.87 37.13 58.46 39.76 72.96 76.28
0 55.61 81.83 61.60 41.21 61.39 44.99 37.11 58.41 39.92 71.61 75.38
Heterogeneous pre-training and fine-tuning decoders: All the previous fine-tuning results were obtained using the pre-trained encoder of a DeCon adaptation, where the pre-training decoder matched the fine-tuning decoder. However, Tab. 5 shows that even if we pre-trained a DeCon adaptation of SlotCon with an FCN decoder and transferred the encoder only for an FPN-based downstream task, we obtained better performance than a baseline SlotCon pre-training. It is also true for the opposite scenario, when FPN decoder was used for pre-training DeCon-ML, and FCN decoder was used for fine-tuning.
Table 5: Heterogeneous pre-training and fine-tuning decoders. We pre-trained a DeCon-SL adaptation of SlotCon with an FCN decoder and transferred the pre-trained encoder to FPN-based models for downstream tasks. Similarly, we pre-trained a DeCon-ML adaptation of SlotCon with an FPN decoder and transferred the pre-trained encoder to FCN-based models for downstream tasks. We averaged fine-tuning results over three runs.
Pretraining $\alpha$ Drop Loss Obj. Det. Obj. Det. Inst. Seg. Semantic Seg.
VOC COCO COCO VOC City
$L_{enc}$ $L_{dec}$ $L_{dds}$ AP AP AP mIoU mIoU
SlotCon 1 0 55.16 40.81 36.80 71.50 75.95
DeCon-ML-FPN 0 0.5 - - - 72.92 76.45
DeCon-SL-FCN 0.5 0 55.29 40.96 36.98 - -
Comparison with state-of-the-art: Table 6 shows that our top-performing DeCon-ML and DeCon-SL adaptations achieved better performance than state-of-the-art methods on COCO object detection and instance segmentation, and VOC semantic segmentation tasks. A more detailed comparison is available in Supplementary Material Section 2.
[…]
Method Obj. Det. Inst. Seg. Semantic Seg.
VOC COCO COCO VOC City
APb APb APm mIoU mIoU
SlotCon ‡ [46] 55.16 40.81 36.80 71.50 75.95
PixCon-SR † [37] 57.55 40.81 36.84 72.95 76.62
DeCon-ML-SlotCon-FPN [ours] 55.70 41.18 37.12 72.92 76.45
DeCon-SL-SlotCon-FCN [ours] - - - 73.01 76.21
DeCon-SL performance in out-of-domain tasks: Table 7 depicts the performance of a model pre-trained on COCO to out-of-domain datasets. We observed a consistent improvement of the DeCon-SL pre-training adaptation compared to the encoder-only pre-training for both REFUGE and ISIC; moreover, this gain increased when the dataset size decreased. Further improvement was achieved when transferring the pre-trained decoder along with the encoder for ISIC dataset. Additionally, pre-training only on ISIC dataset consistently resulted in better performance compared to pre-training on COCO dataset. However, pre-training on REFUGE dataset did not outperform COCO pre-trained approaches, potentially due to the lack of available images (only 400) to develop a strong enough representation during pre-training. In both cases, DeCon-SL pre-training adaptation outperformed encoder-only pre-training, even if we transferred only the encoder. We also observed that pre-trained models outperformed models trained with randomly initialized weights. This gap was more notable in REFUGE dataset (over 20%).
Table 7: Transfer performance on out-of-domain datasets using randomly initialized or pre-trained weights. We pre-trained SlotCon and its DeCon-SL adaptation with an FCN decoder ($\alpha=0.5$) on COCO, REFUGE and ISIC training datasets. We fine-tuned on REFUGE and ISIC datasets under various dataset sizes. We averaged the fine-tuning mIoU results over three runs.
SSL strategy Loss Transfer Pre-trained on COCO Pre-trained on Downstream Data
Refuge ISIC Refuge ISIC
$L_{enc}$ $L_{dec}$ enc dec 5% 25% 100% 5% 25% 100% 100% 100%
Random init. - - - - 49.53 41.41 62.82 75.66 78.38 80.65 62.82 80.65
SlotCon 69.60 77.80 82.83 74.95 79.05 82.52 79.86 83.19
DeCon-SL 71.75 78.92 83.57 75.62 79.56 82.84 80.71 83.66
72.09 77.27 83.25 76.00 79.97 82.94 81.01 83.25
5 Discussion
In this paper, we introduced a contrastive framework adaptation by incorporating a decoder contrastive loss to an existing framework. We showed that our proposed SSL adaptation outperforms previous state-of-the-art methods on COCO object detection and instance segmentation and achieves comparable results on Pascal VOC semantic segmentation (Tab. 6). This adaptation enhances the representation of the pre-trained encoder, improving its transfer performance for dense downstream tasks, regardless of the homogeneity between pre-training and fine-tuning decoders (as shown in Tab. 5). We showed that pre-training an FPN DeCon adaptation of SlotCon leads to better results when using deep supervision (DeCon-ML) compared to the baselines as well as DeCon-SL, both when transferring encoder only and when transferring encoder and decoder. Framework specific hyperparameter tuning of the auxiliary layers could increase further the performance, as shown in Supplementary material Section 3.
We introduced a combined loss where encoder and decoder loss contribute to the total loss in a weighted manner. We observed that increasing the weight for the decoder loss compared to that of the encoder enhances the performance of the pre-trained models on downstream tasks. This could be attributed to the non-competing nature of an encoder and decoder loss compared to approaches that utilize a construction loss for a decoder and a different loss for the encoder in an encoder-decoder architecture [28]. Moreover, we noticed that the encoder loss is optional when pre-training with DeCon-ML, demonstrating the efficiency of the decoder deep supervision and channel dropout combination to ensure a comprehensive pre-training of both encoder and decoder. The lack of skip-connections in DeCon-SL with FCN decoder makes the combined encoder-decoder loss preferable to ensure the best pre-training. We also observed in Tab. 3 that dropout does not provide the same benefits in DeCon-SL compared to DeCon-ML. Hence, channel dropout should only be applied to the different encoder levels when pre-training with a skip-connected decoder and decoder deep supervision (DeCon-ML).
As shown in Tab. 1, DeCon-SL adaptation of an SSL framework also allows for pre-training a decoder efficiently. However, transferring the FPN decoder along the encoder for VOC object detection did not increase detection performance compared to transferring only the encoder, potentially indicating a form of over-fitting of the FPN decoder to the pre-training dataset. This is alleviated by deep supervision in DeCon-ML adaptation as shown in Tab. 3. We showed that DeCon can be used to adapt two different SSL frameworks using different concepts to learn local information (pixel level VS object/group level). This suggests that the proposed DeCon adaptation can be used in any SSL contrastive method designed for dense downstream tasks.
While DeCon-SL adaptation with an FCN decoder benefits DenseCL framework, the added value of the FPN decoder transfer in DeCon-SL adaptation of DenseCL was not obvious on object detection and instance segmentation downstream tasks. However, the added value of the FPN decoder transfer in DeCon-SL adaptation of SlotCon was present across all scenarios, enhancing model performance, possibly due to the region-based contrastive mechanism of SlotCon. We could not investigate the impact of DeCon-ML adaptation on DenseCL due to its extensive memory requirements, which can be attributed to its reliance on a large number of negative pairs.
DeCon’s performance on VOC object detection and Cityscapes semantic segmentation is constrained by its base SSL framework, SlotCon, preventing it from surpassing the state-of-the-art (PixCon) for these tasks(Tab. 6). As an adaptation, DeCon’s effectiveness largely depends on the underlying framework. While integrating DeCon with the latest state-of-the-art framework could set a new benchmark across all datasets, we were unable to implement PixCon due to the unavailability of its code [37].
DeCon adaptation of an SSL framework was also advantageous for fine-tuning on small scale datasets and out-of-domain dense tasks compared to the encoder-only baseline framework (see Tab. 7). This benefit was also observed when pre-training directly on the target dataset. In such cases, DeCon is valuable compared to an encoder-only pre-training whether we transfer the pre-trained decoder or not.
One limitation of our approach is its high memory and computational demand. Pre-training a decoder using DeCon can significantly increase the parameter count and training time, which can restrict this approach to certain SSL frameworks. For example, the DeCon-SL adaptation of SlotCon with FPN decoder could not fit in a single 80GB memory GPU. More details on the parameter count are available in Supplementary Material Section 4. Another limitation of our work is that we only considered ResNet-50 backbone; however, we expect achieving superior results when using more efficient backbones such as EfficientNet [42].
Brempong et al. [4] followed a staged approach, first pre-training the encoder with supervision on annotated datasets, then pre-training the decoder with denoising. However, this multi-stage process relies on labeled data. It could also result in suboptimal feature alignment and increased training complexity. Future works could involve extending our proposed SSL adaptation to offer a multi-stage continual pre-training in a contrastive manner.
6 Conclusion
In this paper, we introduced DeCon—a novel contrastive framework adaptation that enhances the performance of self-supervised learning approaches for dense predictions tasks, such as object detection, semantic segmentation, and instance segmentation. We showed that not-only does this adaptation provide a solution for pre-training decoders, but it also improves the encoder’s representation significantly. We proposed two variants of DeCon: DeCon-SL and DeCon-ML. The former introduces a decoder contrastive loss, and the latter extends DeCon-SL by introducing channel dropout and decoder deep supervision to maximize the encoder’s pre-training power. By integrating these enhancements into an existing framework, we demonstrated that DeCon outperforms or matches state-of-the-art methods across multiple downstream dense prediction tasks.