### v2

# Beyond the Encoder: Joint Encoder-Decoder Contrastive Pre-Training Improves Dense Prediction

Abstract
Contrastive learning methods in self-supervised settings have primarily focused on pre-training encoders, while decoders are typically introduced and trained separately for downstream dense prediction tasks. However, this conventional approach overlooks the potential benefits of jointly pre-training both encoder and decoder. In this paper, we propose DeCon, an efficient encoder-decoder self-supervised learning (SSL) framework that supports joint contrastive pre-training. We first extend existing SSL architectures to accommodate diverse decoders and their corresponding contrastive losses. Then, we introduce a weighted encoder-decoder contrastive loss with non-competing objectives to enable the joint pre-training of encoder-decoder architectures. By adapting an established contrastive SSL framework for dense prediction tasks, DeCon achieves new state-of-the-art results: on COCO object detection and instance segmentation when pre-trained on COCO dataset; across almost all dense downstream benchmark tasks when pre-trained on COCO+ and ImageNet-1K. Our results demonstrate that joint pre-training enhances the representation power of the encoder and improves performance in dense prediction tasks. This gain persists across heterogeneous decoder architectures, various encoder architectures, and in out-of-domain limited-data scenarios.
**footnotetext: Equal contributions for this work. Co-first authors are ordered in alphabetical order.
1 Introduction
The growing demand for efficient Deep Learning (DL) methods stems from their ability to address complex tasks. While traditional approaches like supervised learning can yield high-performing models, they depend on large volumes of high-quality annotated data. Obtaining such annotations is often labor-intensive, expensive, and sometimes impractical, even when abundant data is available. A common workaround is to use pre-trained models, typically encoders trained from scratch on large-scale datasets, to boost performance on downstream tasks with limited labels. Typically, an encoder is pre-trained to capture a representation of the input data, which can then be fine-tuned for downstream tasks such as classification or segmentation. During the downstream fine-tuning phase, randomly initialized layers are typically added on top of the encoder, and the complete architecture is retrained using a supervised learning approach on an annotated dataset. ImageNet [15] is a widely utilized labeled dataset for such pre-training. However, models pre-trained on ImageNet classification task often transfer poorly to downstream dense prediction tasks, such as detection and segmentation [56, 47, 41].
Creating large-scale annotated datasets for every task is impractical. Self-Supervised Learning (SSL) is a promising alternative, enabling models to learn from the large-scale unannotated data to enhance performance on various downstream tasks. These pre-trained models can reach downstream performance close to a fully supervised approach while using significantly fewer labeled examples. Notable approaches such as SimCLR [7], VicReg [2], and MoCo [22] showed promising results by leveraging contrastive learning for downstream classification. However, most SSL frameworks primarily target classification-based downstream tasks. Some efforts, such as PixCon [39] and DenseCL [47], focus on dense prediction tasks like segmentation or object detection. However, these approaches only pre-train encoders for downstream use and rely on local losses adapted from classification-oriented SSL methods without jointly training the decoder.
In this work, we propose a Decoder-aware contrastive learning (DeCon) approach, an encoder-decoder framework that jointly pre-trains encoder and decoder in an SSL manner. DeCon  enhances the representation power of the encoder and prepares it more effectively for integration with decoders in downstream dense prediction tasks. The contributions of this paper are as follows:
  • We propose DeCon-SL, a single-level joint encoder-decoder adaptation of contrastive dense SSL frameworks. We show that jointly pre-training the encoder and decoder enhances the representation power of the encoder and improves downstream dense prediction performance, even when transferring only the pre-trained decoder.
  • We further extend this adaptation by introducing DeCon-ML, which builds on DeCon-SL with a multi-level decoder loss function and an encoder-level channel dropout to promote a comprehensive feature utilization of the encoder. DeCon-ML achieves new state-of-the-art results in COCO object detection and instance segmentation when pre-trained on COCO dataset; and across almost all dense downstream benchmark tasks when pre-trained on COCO+ and ImageNet-1K.
  • We demonstrate that DeCon improves the encoder performance in downstream tasks compared to encoder-only frameworks, even when the decoders used for pre-training and fine-tuning differ.
  • We show that DeCon can achieve gains without increasing parameter count or GPU requirements relative to the original framework.
  • We demonstrate that DeCon remains effective in out-of-domain downstream tasks and yield even greater benefits in scenarios with limited annotated data for fine-tuning.
2 Related Works
Self-Supervised pre-training: In early SSL works, deep learning models were trained on pretext tasks where supervision was derived directly from the data itself. This enabled the models to learn meaningful representations without relying on manual annotations. These tasks included predicting missing parts of an image [40], context prediction [16], solving jigsaw puzzles [36], colorizing grayscale images [54], and predicting rotations [20]. These methods aimed to create surrogate objectives that encouraged models to develop feature representations transferable to downstream tasks. However, they often produced task-specific representations with limited generalizability.
Recent SSL methods prioritize generalizability and move away from hand-crafted pretext tasks. These approaches can be broadly categorized into generative and contrastive methods. Generative approaches aim to learn meaningful representations by modeling the underlying data distribution. This is typically achieved through reconstruction-based methods, where the model predicts corrupted parts of the input [23, 1, 17, 53], or through adversarial methods, where a generator model learns to synthesize realistic samples to challenge a discriminator, thereby encouraging the encoder to extract informative semantic features [50, 26]. Contrastive methods, which are the focus of this research, learn data representations by maximizing the similarity between features extracted from two augmented views of the same image. To avoid model collapse—a phenomenon where the model maps all inputs to the same representation—strategies such as negative sampling [22, 8], stop-gradient mechanisms [9], variance regularization [2, 3], momentum encoders [22], and other architectural and optimization techniques have been employed [5].
Dense pre-training: Although most of the aforementioned frameworks can be used for downstream segmentation or detection, they are not designed specifically for pixel-level—i.e., dense prediction downstream tasks such as instance and semantic segmentation. Consequently, developing SSL approaches that generalize well to such downstream tasks remains an active area of research [48, 47, 46, 3, 37]. The main idea behind these approaches is to shift from a global image similarity to a local similarity. They aim to enforce local similarity based on the location of pixels in an image [52, 46, 37], local pixel features [47], or both [29, 3, 39]. Local similarity can also be enforced in a region-based manner [19, 24, 51, 48]. These methods outperform those relying solely on image-level global similarity when evaluated on dense prediction downstream tasks.
Encoder-only versus encoder-decoder pre-training learning: Encoder-decoder architectures with skip connections between the encoder and decoder layers (i.e., the U-Net family of models) have achieved state-of-the-art results in many segmentation tasks [43, 27, 55]. A common approach for pre-training such models in a self-supervised manner is image reconstruction [30, 4, 6]. However, the presence of skip connections allows a substantial amount of information to bypass the encoder’s final layer—often referred to as the bottleneck layer–resulting in lower-quality embeddings generated by the encoder. Our proposed approach preserves skip connections while addressing the issue of inferior representations generated by the decoder.
Encoder-only contrastive learning frameworks are built solely around an encoder, incorporating components such as predictors and projectors. However, they do not include a decoder. After pre-training, a randomly initialized decoder is attached to the pre-trained encoder in dense prediction downstream tasks to form a complete architecture. While some generative SSL frameworks use a decoder for pre-training [14, 40, 23], to the best of our knowledge, no work has explored pre-training vision models with both encoder and decoder loss in a contrastive learning framework to learn informative data representations.
Figure 1: A: DeCon-SL. Instead of a classical encoder-only pre-training, a decoder is pre-trained alongside the encoder. We mirror the encoder loss at the decoder level and optimize the architecture using a weighted sum of encoder and decoder losses. B: DeCon-ML. Instead of computing the decoder loss at a single level, it is calculated across multiple levels (four in this figure). Additionally, a channel-wise dropout is applied at the output of each encoder level before it is passed through the skip connection to the decoder.
3 Methods
We introduce two architectural adaptations, as depicted in Fig. 1: (A) a single-level decoder loss (DeCon-SL) and (B) a multi-level decoder loss (DeCon-ML).
3.1 DeCon-SL
Given a vision SSL framework that pre-trains an encoder using a teacher-student architecture [25], we retain the encoders and any existing “auxiliary layers” (e.g., predictors, projection heads), and add decoders and their corresponding auxiliary layers for both teacher and student networks. For all models, the auxiliary layers are designed to match those of the encoder, with adjustments applied to align with the decoder’s output feature size. We obtain two losses, one is computed from the encoder features while the other one is computed from the decoder features. Consequently, the loss function is defined as follows:
$Loss=\alpha\times L_{enc}+(1-\alpha)\times L_{dec}$ (1)
where, $\alpha$ represents the contribution of the encoder loss in the objective function.
The DeCon-SL architecture is illustrated in Fig. 1 A, where $f_{\theta}$ and $f_{\epsilon}$ represent student and teacher encoders, respectively. $aux^{enc}_{\theta}$ and $aux^{enc}_{\epsilon}$ are the SSL framework’s auxiliary layers of the student and teacher encoders, respectively. Auxiliary layers can include any projector, predictor heads, or any layers used in the loss definition of the original SSL framework. $g_{\theta}$ and $g_{\epsilon}$ are the decoders of the student and teacher networks, and $aux^{dec}_{\theta}$ and $aux^{dec}_{\epsilon}$ their respective auxiliary layers. $g_{\theta}$ and $g_{\epsilon}$ take the output of encoder $f_{\theta}$ and $f_{\epsilon}$ as input, respectively. Auxiliary layers $aux^{enc}_{\theta}$ and $aux^{enc}_{\epsilon}$ are responsible for generating a representation from the encoder output, that is used to compute a contrastive loss for the encoders $L^{encoder}$. The contrastive loss, $L^{decoder}$ is calculated from the representation obtained from the auxiliary layers of the decoders, i.e., $aux^{dec}_{\theta}$ and $aux^{dec}_{\epsilon}$. $f_{\epsilon}$ and $g_{\epsilon}$ are updated using the exponential moving average (EMA) of $f_{\theta}$ and $g_{\theta}$, respectively.
3.2 DeCon-ML
To promote a comprehensive usage of the encoder’s parameters, we introduce a multi-layer contrastive decoder loss. This updated framework consists of a single loss for the encoder and multiple loss components for the decoder in a deep supervision manner along with channel dropout, which we describe in detail below. Figure 1 B illustrates the DeCon-ML adaptation of a given SSL framework.
Channel dropout: The dropout layer is typically used to prevent overfitting, improve generalization, and act as an implicit ensemble [44, 45]. However, we extend its use by applying dropout to channels transferred through skip connections between the encoder and decoder. In DeCon-ML, we apply channel dropout to the output layers of the different encoder levels by zero-ing out entire channels of the feature maps when passing these outputs to the decoder. No channel is zeroed-out when the inputs pass through the encoder. This channel dropout prevents the model from over-relying on specific features shared through skip connections. It encourages a more comprehensive use of the encoder’s parameters at the different levels, leading to a richer and more powerful data representation learned, while still preserving the input information content as it passes through the encoder to its bottleneck.
Decoder deep supervision: An encoder-decoder SSL framework allows us to retrieve meaningful features from multiple levels of the decoder. We can then pre-train the architecture with multiple losses using decoder deep supervision. If the decoder is connected to the encoder at multiple levels through lateral/skip connections, decoder deep supervision strengthens the representation power of the encoder at different levels [42]. Since a decoder passes the encoder bottleneck representation in a bottom-up manner to its upper level layers, the final encoder bottleneck representation remains a non-negligible objective of the training since it impacts all decoder level losses. In contrast, applying deep supervision at multiple encoder stages in an encoder-only SSL framework might prevent the creation of a strong semantic representation at the encoder bottleneck. To perform the decoder deep supervision, we add auxiliary layers to each decoder level, and a loss is computed at each level between the teacher and student representations. When performing decoder deep supervision, the encoder loss will remain the same as Eq. 1, but the decoder loss will be updated as follows:
$Loss=$ $\alpha\times L_{enc}+(1-\alpha)\times L_{dds}$ (2)
$L_{dds}=$ $\frac{1}{j}\times\sum_{i=1}^{j}L_{dec_{i}}$ (3)
where $j$ is the number of decoder levels at which a loss will be calculated and $L_{dec_{i}}$ is the loss computed at level $i$. The final decoder loss in Eq. 2 is thus an average of the losses at the different levels.
3.3 Implementation details
Architecture: We use a ResNet-50 encoder [21] as a backbone and implement a Fully Convolutional Network (FCN) [34] and a skip-connected Feature Pyramid Network (FPN) [32] as two decoder architectures in the DeCon framework. We follow the mmseg [11] FCN implementation used in SlotCon [48] and customize the FPN implementation from Detectron2 [49] version 0.6. Details on these decoder architectures are provided in Supplementary Material Section 1. We use DeCon to adapt the SlotCon [48] framework. Additional experiments are conducted using DenseCL [47] and PixPro [52] framework or a ConvNeXt-small [33] backbone.
For SlotCon, the student and teacher encoders, and every auxiliary layers—i.e., the respective projectors and semantic groupings, the slot predictor and the prototypes—are kept unchanged. Two decoders, a student and a teacher decoder (either FCN or FPN), that take the output of their corresponding encoder as input are added to the framework. The auxiliary layers are replicated at the decoder level. In the case of decoder deep supervision with FPN decoders (DeCon-ML), auxiliary layers are replicated independently at the four decoder levels. No deep supervision is performed in the DeCon adaptation of SlotCon with an FCN decoder (DeCon-SL). The decoder projectors differ from the encoder projectors only in input channels: 2048 for the encoder and 256 for the FPN and FCN decoders. We follow SlotCon by updating the student branch and the corresponding encoder and decoder slot predictors with Backpropagation, the teacher branch with an EMA of the first one, and the prototypes centers with an EMA of the previous batch centers seen during training. Our implementation is based on the official SlotCon repository.
We adapt in a similar way DenseCL and PixPro into a DeCon-SL framework using an FCN decoder. More details are given in Supplementary Material 2. We adapt DenseCL implementation from mmselfsup package [12] version 1.0.0 and PixPro from the original implementation.
Although the backbone pre-trained remains unchanged, adding multiple decoder levels and associated loss functions increases the overall number of parameters in the DeCon-ML architecture. To enable a fair comparison with the original SlotCon framework using a similar parameter budget, we implemented a reduced version of DeCon-ML that includes only the first two decoder levels originating from the encoder bottleneck. Additionally, the hidden dimension in the decoder projector was reduced from 4096 to 2048. This smaller version, which matches the parameter count of the original SlotCon framework (when $\alpha=0$), is referred to as DeCon-ML-S, while the full model with four decoder levels and unmodified configuration is denoted DeCon-ML-L. Detailed figures of DeCon-SL, DeCon-ML-S and DeCon-ML-L adaptations of SlotCon are available in the Supplementary Material Section 3.
Pre-training setup: We pre-train DeCon on COCO 2017 dataset [31], COCO+ (train2017 + unlabeled2017 datasets) and ImageNet-1K datasets [15]. The augmentations and hyperparameters—learning rate, weight decay, schedulers, input shape, momentum, and so on—are kept the same as used in the original SSL framework. Models are pre-trained on COCO and COCO+ for 800 epochs, and for 200 epochs on ImageNet. The SlotCon ResNet-50 models are pre-trained using a LARS optimizer, with a total batch size of 512 across all available GPUs. The base learning rate is 1.0 scaled with the batch size and updated with a cosine learning rate decay schedule using weight decay of $10^{-5}$ and a warm-up period of 5 epochs. The ConvNeXt-small backbone is pre-trained in SlotCon and DeCon-ML using an AdamW optimizer [35], a base learning rate of 0.002 and weight-decay 0.05. Remaining hyperparameters are kept the same as for the ResNet-50 implementation. The pre-trainings using FPN decoders are done using eight NVIDIA H100 GPUs (80GB each). The pre-trainings without decoder and with FCN decoders are conducted on one 80GB NVIDIA A100 GPU.
For DenseCL, DeCon-SL is pre-trained using an SGD optimizer with a total batch size of 256. The base learning rate is 0.3 and is updated using a cosine learning rate decay schedule (0.0001 weight decay, 0.9 momentum). The pre-training experiments are conducted with eight NVIDIA A100 GPUs (40GB each). For PixPro, DeCon-SL is pre-trained using a LARS optimizer for 800 epochs with a batch size of 512 and a learning rate of 1.0 with a weight decay of 1e-5. The pre-trainings are conducted with eight NVIDIA A100 GPUs (40GB each).
To assess the DeCon framework on non scene-centric data, we pre-train SlotCon and a DeCon-SL adaptation of SlotCon with an FCN decoder on two small-scale datasets: ISIC 2017 and REFUGE described in Sec. 3.4. For ISIC dataset, we pre-train for 800 epochs with a batch size of 256 and a base learning rate of 1.0 linearly scaled with the batch size. For REFUGE, we pre-train for 2400 epochs with a batch size of 192 and the same learning rate as for ISIC. These pre-trainings are performed with one NVIDIA A6000 GPU with 48GB of memory.
3.4 Evaluation protocol
We evaluate the different pre-trained models on object detection, semantic segmentation, and instance segmentation as downstream tasks. Unless stated otherwise, we only transfer the pre-trained encoder for the fine-tuning experiments. We initialize the network with a pre-trained backbone and fine-tune end-to-end in all cases. All fine-tuning results reported in this paper are averaged over three independent runs. The pre-trainings, however, are only run once. All fine-tuning experiments related to the same dataset are run on a similar platform with the same hardware. We perform object detection and instance segmentation using Detectron2 package [49] version 0.6, whereas mmsegmentation [11] version 0.30.0 is used for semantic segmentation tasks.
Object detection and instance segmentation: We perform object detection and instance segmentation on COCO dataset [31]. We fine-tune a Mask R-CNN with FPN backbone following SlotCon [48].
Semantic segmentation: We evaluate the semantic segmentation performance of a pre-trained ResNet-50 encoder with an FCN decoder on Pascal VOC [18] and Cityscapes [13] following SlotCon [48] and PixCon [39].
Semantic segmentation on out-of-domain datasets: We assess generalization on REFUGE [38] and ISIC 2017 [10] datasets using the same ResNet-50-FCN architecture, fine-tuned on 5, 25, and 100% of the training data.
More details on the training setups and hyperparameters are provided in the Supplementary Material Section 2.
4 Results
4.1 Experiments
Table 1: Performance of DeCon-SL ($\alpha=0.25$) and DeCon-ML Small and Large ($\alpha=0$ and $dropout=0.5$) adaptations of SlotCon trained with a ResNet-50 backbone. The architecture was pre-trained for 800 epochs on COCO and COCO+ and for 200 epochs on ImageNet-1K, and then fine-tuned on various segmentation and object detection tasks. Only the pre-trained encoder was transferred for the fine-tuning. Results for PixCon pre-trained on ImageNet-1K are not reported, as the corresponding checkpoint or evaluation was not made available in the original publication. All fine-tuning results were averaged over three runs. means that we reimplemented the framework, * means that the results are taken from the original paper.
Pret. Dataset Framework Pret. Dec. Object Detection Instance Segmentation Semantic Segmentation
COCO COCO VOC City
AP AP50 AP75 AP AP50 AP75 mIoU mIoU
COCO PixCon [39]* - 40.81 60.97 44.80 36.80 57.93 39.62 72.95 76.62
SlotCon [48] - 40.81 60.95 44.37 36.80 57.98 39.54 71.50 75.95
DeCon-SL FCN 40.97 61.22 44.81 36.92 58.12 39.78 73.01 76.21
DeCon-ML-S FPN 40.97 61.20 44.71 36.94 58.20 39.63 72.80 76.21
DeCon-ML-L FPN 41.18 61.38 44.91 37.12 58.35 39.94 72.92 76.45
COCO+ PixCon [39]* - 41.2 - - 37.1 - - 73.9 77.0
SlotCon [48] - 41.63 62.10 45.67 37.57 59.07 40.45 73.93 76.43
DeCon-SL FCN 41.86 62.43 45.73 37.75 59.40 40.48 74.46 76.65
DeCon-ML-L FPN 42.08 62.42 46.13 37.84 59.41 40.75 75.36 77.00
ImageNet-1K SlotCon [48] - 41.69 62.07 45.59 37.59 58.97 40.49 75.02 76.15
DeCon-ML-L FPN 41.80 62.12 45.73 37.73 59.08 40.68 75.40 76.51
ResNet-50 backbone: Table 1 presents the results of fine-tuning a ResNet-50 backbone pre-trained with the two best pre-training frameworks for dense downstream tasks and DeCon. A more comprehensive comparison with additional previously published methods is provided in the Supplementary Material 4. DeCon-ML-L is consistently establishing a new SATO for all tasks, except when pre-training is performed on COCO dataset for Pascal VOC and Cityscapes semantic segmentation fine-tuning, where it reduces the gap between PixCon and SlotCon. This supports the fact that the DeCon adpatation is robust across datasets. Secondly, we can see that all of DeCon versions consistently improve upon the base SlotCon framework for any of the pre-training datasets used. Notably, DeCon-ML-S, whose parameter count is similar to SlotCon, exhibits better performance than SlotCon.
Table 2: Performance of DeCon-ML-L ($\alpha=0$ and $dropout=0.5$) adaptation of SlotCon using a ConvNeXt-small encoder. The architecture was pre-trained for 800 epochs on COCO, and then fine-tuned on various segmentation and object detection tasks. All fine-tuning results were averaged over three runs.
Pret. Dataset Framework Pret. Dec. Obj. Det. Inst. Seg. Sem. Seg.
COCO COCO VOC
AP AP mIoU
COCO SlotCon - 44.07 39.67 73.24
DeCon-ML-L FPN 44.71 40.37 73.81
ConvNeXt backbone: Table 2 presents the results of the fine-tuning of models pre-trained with a ConvNeXt-small backbone using DeCon-ML-L and the original SlotCon. A consistent improvement is seen with our encoder-decoder framework, confirming the robustness of our method to more modern and bigger backbones. Notably, the performance of the model as well as the performance improvement appear larger than for a smaller ResNet-50 backbone.
Table 3: Performance of DeCon-SL ($\alpha=0.5$) adaptation of DenseCL and PixPro using a ResNet-50 encoder. The architectures were pre-trained for 800 epochs on COCO, and then fine-tuned on Pascal VOC and Cityscapes semantic segmentation. All fine-tuning results were averaged over three runs.
Pret. Dataset Framework Pret. Dec. Transfer Sem. Seg.
VOC City
mIoU mIoU
COCO DenseCL - Enc 71.11 75.87
DeCon-SL (DenseCL) FCN Enc 71.31 76.04
Enc + Dec 71.59 76.17
PixPro - Enc 72.13 75.84
DeCon-SL-(PixPro) FCN Enc 72.18 76.16
Enc + Dec 73.08 76.19
Generalizability across frameworks: Table 3 presents semantic segmentation fine-tuning results of models pre-trained with DenseCL and PixPro, and their DeCon-SL adaptation. The performance of the model pre-trained with a decoder is improved compared to the base framework and improves even further when the pre-trained decoder is transferred, supporting the hypothesis that DeCon could be adapted to multiple frameworks. Transferring the pre-trained decoder however, is not always beneficial to downstream performance in the case of SlotCon. More details can be found in Supplementary Material 5.
Out-of-domain tasks: Table 4 depicts the fine-tuning performance of a model pre-trained on COCO to out-of-domain datasets. We observed a consistent improvement of the DeCon-SL pre-training adaptation compared to the encoder-only pre-training for both REFUGE and ISIC; moreover, this gain increased when the dataset size decreased. Further improvement was achieved when transferring the pre-trained decoder along with the encoder for ISIC dataset. Additionally, pre-training only on ISIC dataset consistently resulted in better performance compared to pre-training on COCO dataset. However, pre-training on REFUGE dataset did not outperform COCO pre-trained approaches, potentially due to the lack of available images (only 400) to develop a strong enough representation during pre-training. In both cases, DeCon-SL pre-training adaptation outperformed encoder-only pre-training, even if only the encoder was transferred. We also observed that pre-trained models outperformed models trained with randomly initialized weights. This gap was more notable in REFUGE dataset (over 20%).
Table 4: Transfer performance on out-of-domain datasets using randomly initialized or pre-trained weights. We pre-trained SlotCon and its DeCon-SL adaptation with a ResNet-50 encoder and an FCN decoder ($\alpha=0.5$) on COCO, REFUGE and ISIC training datasets. We fine-tuned on REFUGE and ISIC datasets under various dataset sizes. We averaged the fine-tuning mIoU results over three runs.
SSL strategy Loss Transfer Pre-trained on COCO Pre-trained on Downstream Data
REFUGE ISIC REFUGE ISIC
$L_{enc}$ $L_{dec}$ enc dec 5% 25% 100% 5% 25% 100% 100% 100%
Random init. - - - - 49.53 41.41 62.82 75.66 78.38 80.65 62.82 80.65
SlotCon 69.60 77.80 82.83 74.95 79.05 82.52 79.86 83.19
DeCon-SL 71.75 78.92 83.57 75.62 79.56 82.84 80.71 83.66
72.09 77.27 83.25 76.00 79.97 82.94 81.01 83.25
4.2 Ablation studies
Table 5 presents an ablation study of the components in our DeCon-ML-L framework. The results show that while pre-training with a decoder and deep-supervision provides a modest benefit to downstream fine-tuning performance on COCO, even when only the encoder is transferred, the main performance improvement arises from the use of channel dropout in conjunction with decoder deep supervision. This combination yields the most substantial gains, highlighting the critical role of channel dropout when paired with deep supervision in our approach.
Table 5: Ablation of the different components of DeCon-ML-L and their impact on the downstream performance. We averaged all fine-tuning results over three runs. $\Delta$ is the performance gap between the previous framework and the framework updated with the corresponding new component.
Framework Improvement COCO Obj Det. COCO Inst. Seg.
AP $\Delta$ AP $\Delta$
SlotCon 40.81 - 36.80 -
+FPN Decoder (Multi Level Loss) 40.85 +0.04 36.83 +0.03
+Dropout = 0.5 41.10 +0.25 37.05 +0.22
+Enc-Dec Loss Weight ($\alpha$) Tuning (DeCon-ML-L) 41.18 +0.08 37.12 +0.07
Table 6 shows DeCon-ML performance with different values of encoder loss weight $\alpha$ and channel dropout. A channel dropout probability of 0.5 provided the best downstream performance for DeCon-ML-L. Hence, this value was chosen to ablate $\alpha$. An $\alpha$ value of 0 gave the best downstream performance meaning that the decoder loss can replace the encoder one and pre-train efficiently an encoder on its own. For DeCon-SL, an $\alpha$ value of 0.25 resulted in the best downstream performance, suggesting that having an encoder loss at the bottleneck is still beneficial to the pre-training in the absence of skip connections.
Table 6: Impact of the channel dropout probability and of the encoder loss weight $\alpha$ on DeCon. We pre-trained DeCon adaptations of SlotCon framework on COCO dataset using different parameters and evaluate the pre-trained encoder fine-tuning on downstream tasks that use the same decoder architecture. We averaged all fine-tuning results over three runs.
Abla tion Frame work Deco der $\alpha$ Drop. Obj. Det. Inst. Seg. Frame work Deco der $\alpha$ Sem. Seg.
COCO COCO VOC City
AP AP mIoU mIoU
Dropout
DeCon-ML-L
FPN
0.50 0 40.85 36.83 - - - - -
0.25 40.95 36.97 - - - - -
0.50 41.10 37.05 - - - - -
0.75 41.02 37.05 - - - - -
$\alpha$
0.75 0.50 40.98 36.90
DeCon-SL (No Drop)
FCN
0.75 71.98 76.18
0.50 41.10 37.05 0.50 72.42 75.62
0.25 41.05 37.13 0.25 73.01 76.21
0 41.18 37.11 0 72.17 75.71
5 Discussion
Table 7: COCO and COCO+ pretraining downstream performance improvement with DeCon-ML-L and Cohen’s d value. PixCon performance was extracted from the original paper.
Method COCO obj det. COCO Inst. Seg. VOC Sem. Seg. City Sem. Seg.
$\Delta$AP $\Delta$AP $\Delta$mIoU $\Delta$mIoU
COCO Pre-training Performance Improvement $\Delta$ (calculated from the respective papers)
SlotCon (2022) (Comp. to prev. works) +0.50 +0.40 -0.40 +0.10
PixCon (2024) (Comp. to SlotCon) 0.00 +0.06 +1.30 +0.51
DeCon-ML-L (Comp. to SlotCon) +0.37 +0.32 +1.42 +0.50
Cohen’s d (DeCon-ML-L VS SlotCon) 2.35 1.94 6.37 2.51
COCO+ Pre-training Performance Improvement $\Delta$ (calculated from the respective papers)
SlotCon (2022) (Comp. to prev. works) +0.60 +0.50 Not avail. Not avail.
PixCon (2024) (Comp. to SlotCon) -0.50 -0.50 -0.20 +0.40
DeCon-ML-L (Comp. to SlotCon) +0.45 +0.27 +1.43 +0.57
Cohen’s d (DeCon-ML-L VS SlotCon) 5.42 3.21 5.25 1.22
In this paper, we introduced a contrastive framework adaptation by incorporating a decoder contrastive loss to an existing framework. We showed that our proposed SSL adaptation outperforms previous state-of-the-art methods on COCO object detection and instance segmentation and achieves comparable results on Pascal VOC semantic segmentation when pre-trained on COCO dataset, and consistently outperforms previous methods when pre-trained on COCO+, and ImageNet except for Cityscapes semantic segmentation (Tab. 1). This adaptation enhances the representation of the pre-trained encoder, improving its transfer performance for dense downstream tasks, regardless of the homogeneity between pre-training and fine-tuning decoders (as shown in Tab. 1).
We introduced a combined loss where encoder and decoder losses contribute to the total loss in a weighted manner.
Figure 2: Pre-training loss dynamics for DeCon-ML-L and SlotCon.
Figure 2 illustrates the loss behavior across different pre-trainings when both the encoder and decoder are used. As shown, decoder pre-training does not influence the encoder’s loss dynamics, supporting the fact that the encoder and decoder loss terms are conceptually non-competing. We also observed that increasing the weight for the decoder loss compared to that of the encoder enhances the performance of the pre-trained models on downstream tasks. This could be attributed to the non-competing nature of an encoder and decoder loss compared to approaches that utilize a reconstruction loss for a decoder and a different loss for the encoder in an encoder-decoder architecture [28]. Additionally, Fig. 3 presents the slots learned by SlotCon encoder and by the different decoder levels of DeCon-ML. It demonstrates that the decoder learns features that capture similar concepts as the encoder’s, while being more spatially precise.
Figure 3: Slots, as defined in SlotCon, learned at different outputs of our architecture. The top left image from COCO validation dataset is overlayed with the slot that was the most represented in original SlotCon’s encoder output feature map, resized to the image’s shape. Other slots displayed are the slots that had the biggest overlap with SlotCon’s main encoder output slot. More details on the slots creation are available in Supplementary Material 6.
Moreover, we noticed that the encoder loss is optional when pre-training with DeCon-ML, demonstrating the efficiency of the decoder deep supervision and channel dropout combination to ensure a comprehensive pre-training of both encoder and decoder. The lack of skip-connections in DeCon-SL with FCN decoder makes the combined encoder-decoder loss preferable to ensure the best pre-training. In Supplementary Material Section 7, we provide experiments on using DeCon-SL with an FPN decoder. We observed that with an FPN decoder, dropout does not provide the same benefits in DeCon-SL compared to DeCon-ML. Hence, channel dropout should only be applied to the different encoder levels when pre-training with a skip-connected decoder and decoder deep supervision (DeCon-ML). Framework specific hyperparameter tuning of the decoder auxiliary layers could also increase the fine-tuning performance (shown in Supplementary Material Section 8).
We demonstrated that our method is effective with a different type of encoders. Specifically, pre-training a ConvNeXt-small backbone, which is both more efficient and larger than the ResNet-50 backbone, also resulted in superior downstream performance when integrated with a decoder in the DeCon-ML-L framework. These results attest to the scalability and practical utility of the proposed approach for pre-training modern backbone architectures.
DeCon was also used to successfully adapt three different SSL frameworks using different concepts (pixel level VS object/group level) to learn local information (Tab. 3). This suggests that the proposed DeCon adaptation could be used in other SSL contrastive methods designed for dense downstream tasks.
Table 8: Pre-training parameter count, GPU setup, and time required per epoch for the different architectures. We study various pre-training setups on the COCO dataset using the SlotCon framework. Module sizes are given in millions of parameters.
Pre-training Enc. size Dec. size Enc. Aux. Layers Size Dec. Aux. Layers Size Full Archi. Size GPU setup Time per epoch (s)
SlotCon 23.51 x 2 0 71.51 0 118.52 1 * A100 80GB 255
DeCon-SL-FCN ($\alpha$ = 0.25) 23.51 x 2 5.31 x 2 71.51 56.83 185.97 1 * A100 80GB 300
DeCon-ML-S ($\alpha$ = 0) 23.51 x 2 3.35 x 2 0 59.08 110.03 1 * A100 80GB 280
DeCon-ML-L ($\alpha$ = 0) 23.51 x 2 3.35 x 2 0 227.31 281.02 8 * H100 80GB 165
DeCon adaptation of an SSL framework provided notable advantages for fine-tuning on small-scale datasets and out-of-domain dense prediction tasks, outperforming the encoder-only baseline (see Tab. 4). This improvement persisted even when pre-training was conducted directly on the small target datasets. Furthermore, DeCon achieved superior performance compared to related frameworks when only a limited fraction (5–25%) of the original labeled data was available for fine-tuning. These results highlight the value of DeCon in settings with very limited annotated samples or small-scale, domain-specific datasets, such as those in medical or agricultural applications, regardless of whether the pre-trained decoder is transferred.
Performance of deep learning models can vary across multiple runs due to inherent stochasticity. To mitigate the impact of randomness while considering computational resource constraints, each experiment was conducted three times, with all fine-tunings for a specific downstream task performed on similar hardware. The performance improvements achieved with our DeCon framework compared to the original framework are statistically significant, as evidenced by large Cohen’s $d$ values (>0.8>0.8> 0.8) in Tab. 7, exceeding two standard deviations (see Supplementary Material Section 9), and $p$-values less than 0.05 (0.012 for COCO object detection and 0.048 for instance segmentation) from Wilcoxon signed-rank tests. Details on the computation of these values are provided in Supplementary Material 10. Furthermore, the performance gains shown in Tab. 7 are comparable to those reported in prior studies.
One limitation of our approach is its high memory and computational demand. However, we have shown in Tab. 1 that if GPU is a constraint, the user can decide to reduce the number of decoder levels used in the pre-training and tune the decoder projector layers to reduce the parameter count of DeCon-ML architecture and match the parameter count of the original framework (as shown in Tab. 8), while still improving the performance of the pre-training.
Brempong et al. [4] followed a staged approach, first pre-training the encoder with supervision on annotated datasets, then pre-training the decoder with denoising. However, this multi-stage process relies on labeled data. It could also result in suboptimal feature alignment and increased training complexity. Future works could involve extending our proposed DeCon adaptation to offer a multi-stage continual pre-training in a contrastive manner.
6 Conclusion
In this paper, we introduced DeCon—a novel contrastive framework adaptation that enhances the performance of self-supervised learning approaches for dense predictions tasks, such as object detection, semantic segmentation, and instance segmentation. We showed that this adaptation improves the encoder’s representation significantly. We proposed two variants of DeCon: DeCon-SL and DeCon-ML. The former introduces a decoder contrastive loss, and the latter extends DeCon-SL by introducing channel dropout and decoder deep supervision to maximize the encoder’s pre-training power. By integrating the proposed contrastive joint encoder-decoder pre-training strategy into an existing framework, we show that DeCon consistently outperforms or matches state-of-the-art methods across diverse dense prediction tasks, highlighting the effectiveness of this unified pre-training approach.
Acknowledgments
Computing resources were provided by the University of British Columbia Cloud Innovation Centre, powered by Amazon Web Services https://cic.ubc.ca/. This research was also partly enabled by support provided by the Digital Research Alliance of Canada https://alliancecan.ca/, and the Research Computing Services group at the University of Calgary https://rcs.ucalgary.ca/RCS_Home_Page.