### v3

# Beyond the Encoder: Joint Encoder-Decoder Contrastive Pre-Training Improves Dense Prediction

Abstract
Contrastive learning methods in self-supervised settings have primarily focused on pre-training encoders, while decoders are typically introduced and trained separately for downstream dense prediction tasks. However, this conventional approach overlooks the potential benefits of jointly pre-training both encoder and decoder. In this paper, we propose DeCon, an efficient encoder-decoder self-supervised learning (SSL) framework that supports joint contrastive pre-training. We first extend existing SSL architectures to accommodate diverse decoders and their corresponding contrastive losses. Then, we introduce a weighted encoder-decoder contrastive loss with non-competing objectives to enable the joint pre-training of encoder-decoder architectures. By adapting a contrastive SSL framework for dense prediction, DeCon establishes consistent state-of-the-art performance on most of the evaluated tasks when pre-trained on Imagenet-1K, COCO and COCO+. Notably, when pre-training a ResNet-50 encoder on COCO dataset, DeCon improves COCO object detection and instance segmentation compared to the baseline framework by +0.37 AP and +0.32 AP, respectively, and boosts semantic segmentation by +1.42 mIoU on Pascal VOC and by +0.50 mIoU on Cityscapes. These improvements generalize across recent backbones, decoders, datasets, and dense tasks beyond segmentation and object detection, and persist in out-of-domain scenarios, including limited-data settings, demonstrating that joint pre-training significantly enhances representation quality for dense prediction. Code is available at https://github.com/sebquetin/DeCon.git.
**footnotetext: Equal contributions. Co-first authors are ordered in alphabetical order. Corresponding author: T. Ghosh (tapotosh.ghosh@ucalgary.ca)
1 Introduction
The growing demand for efficient Deep Learning (DL) methods stems from their ability to address complex tasks. While traditional approaches like supervised learning can train high-performing models, they depend on large volumes of high-quality annotated data. Obtaining such annotations is often labor-intensive, expensive, and sometimes impractical, even when abundant data is available. A common workaround is to use pre-trained models, typically encoders trained from scratch on large-scale datasets, to boost performance on downstream tasks with limited labels. Typically, an encoder is pre-trained to capture a representation of the input data, which can then be fine-tuned for downstream tasks such as classification or segmentation. During the downstream fine-tuning phase, randomly initialized layers are typically added on top of the encoder, and the complete architecture is retrained using a supervised learning approach on an annotated dataset. ImageNet [18] is a widely utilized labeled dataset for such pre-training. However, models pre-trained on ImageNet classification task often transfer poorly to downstream dense prediction tasks, such as detection and segmentation [65, 53, 45].
Creating large-scale annotated datasets for every task is impractical. Self-Supervised Learning (SSL) is a promising alternative, enabling models to learn from the large-scale unannotated data to enhance performance on various downstream tasks. These pre-trained models can reach downstream performance close to a fully supervised approach while using significantly fewer labeled examples. Notable approaches such as SimCLR [9], VicReg [3], and MoCo [25] showed promising results by leveraging contrastive learning for downstream classification. However, most SSL frameworks primarily target classification-based downstream tasks. Some efforts, such as PixCon [42] and DenseCL [53], focus on dense prediction tasks like segmentation or object detection. However, these approaches only pre-train encoders for downstream use and rely on local losses adapted from classification-oriented SSL methods without jointly training the decoder.
In this work, we propose a Decoder-aware contrastive learning (DeCon) approach, an encoder-decoder framework that jointly pre-trains encoder and decoder in an SSL manner. DeCon enhances the representation power of the encoder, and prepares it more effectively for integration with decoders in downstream dense prediction tasks. The contributions of this paper are as follows:
  • We propose DeCon-SL, a single-level joint encoder-decoder adaptation of contrastive dense SSL frameworks. We show that jointly pre-training the encoder and decoder enhances the representation power of the encoder and improves downstream dense prediction performance, even when transferring only the pre-trained encoder.
  • We further extend this adaptation by introducing DeCon-ML, which builds on DeCon-SL with a multi-level decoder loss function and an encoder-level channel dropout to promote a comprehensive encoder feature utilization.
  • DeCon achieves new state-of-the-art (SOTA) results in COCO object detection and instance segmentation, and Pascal VOC, Cityscapes and ADE20K semantic segmentation when pre-trained on ImageNet-1K and COCO+.
  • We show that DeCon consistently improves performance across different backbones, a range of dense downstream tasks beyond object detection, semantic and instance segmentation as well as in various out-of-domain scenarios.
  • We show that DeCon can achieve gains without increasing parameter count and with comparable GPU cost relative to the original framework, e.g. SlotCon [55].
2 Related Work
Self-Supervised pre-training: In early SSL works, deep learning models were trained on pretext tasks where supervision was derived directly from the data itself. This enabled the models to learn meaningful representations without relying on manual annotations. These tasks included predicting missing parts of an image [43], context prediction [19], solving jigsaw puzzles [39], colorizing grayscale images [62], and predicting rotations [23]. These methods aimed to create surrogate objectives that encouraged models to develop feature representations transferable to downstream tasks. However, they often produced task-specific representations with limited generalizability.
Recent SSL methods prioritize generalizability and move away from hand-crafted pretext tasks. These approaches can be broadly categorized into generative and contrastive methods. Generative approaches learn representations by modeling the underlying data distribution. This is typically achieved through reconstruction-based methods, where the model predicts corrupted parts of the input [24, 1, 20, 61], or through adversarial methods, where a generator model learns to synthesize realistic samples to challenge a discriminator, thereby encouraging the encoder to extract informative semantic features [57, 29]. Contrastive methods, which are the focus of this research, learn data representations by maximizing the similarity between features extracted from two augmented views of the same image. To avoid model collapse—a phenomenon where the model maps all inputs to the same representation—strategies such as negative sampling [25, 10], stop-gradient mechanisms [11], variance regularization [3, 4], momentum encoders [25], and other architectural and optimization techniques have been employed [6].
Dense pre-training: Although most of the aforementioned frameworks can be used for downstream segmentation or detection, they are not designed specifically for pixel-level—i.e., dense prediction downstream tasks such as instance and semantic segmentation. Consequently, developing SSL approaches that generalize well to such downstream tasks remains an active area of research [55, 53, 52, 4, 40]. The main idea behind these approaches is to shift from a global image similarity to a local similarity. They aim to enforce local similarity based on the location of pixels in an image [60, 52, 40], local pixel features [53], or both [32, 4, 42]. Local similarity can also be enforced in a region-based manner [22, 27, 59, 55]. These methods outperform those relying solely on image-level global similarity when evaluated on dense prediction downstream tasks.
Encoder-only versus encoder-decoder pre-training: Encoder-decoder architectures with skip connections between the encoder and decoder layers (i.e., the U-Net family of models) have achieved state-of-the-art results in many segmentation tasks [48, 30, 64]. A common approach for pre-training such models in a self-supervised manner is image reconstruction [33, 5, 8]. However, the presence of skip connections allows a substantial amount of information to bypass the encoder’s final layer—often referred to as the bottleneck layer–resulting in lower-quality embeddings generated by the encoder. Our proposed approach preserves skip connections while addressing the issue of inferior representations generated by the decoder.
Encoder-only contrastive learning frameworks are built solely around an encoder, incorporating components such as predictors and projectors. However, they do not include a decoder. After pre-training, a randomly initialized decoder is attached to the pre-trained encoder in dense prediction downstream tasks to form a complete architecture. While some generative SSL frameworks use a decoder for pre-training [17, 43, 24], a clear gap remains in exploring vision model pre-training that jointly employs encoder and decoder losses within a unified contrastive learning framework to learn informative data representations.
Figure 1: A: DeCon-SL. Instead of a classical encoder-only pre-training, a decoder is pre-trained alongside the encoder. We mirror the encoder loss at the decoder level and optimize the architecture using a weighted sum of encoder and decoder losses. B: DeCon-ML. Instead of computing the decoder loss at a single level, it is calculated across multiple levels (four in this figure). Additionally, a channel-wise dropout is applied at the output of each encoder level before it is passed through the skip connection to the decoder.
3 Methods
We introduce two architectural adaptations, as depicted in Fig. 1: (A) a single-level decoder loss (DeCon-SL) and (B) a multi-level decoder loss (DeCon-ML).
3.1 DeCon-SL
Given a vision SSL framework that pre-trains an encoder using a teacher-student architecture [28], we retain the encoders and any existing “auxiliary layers” (e.g., predictors, projection heads), and add decoders and their corresponding auxiliary layers for both teacher and student networks. For all models, the decoder auxiliary layers are designed to match those of the encoder, with adjustments applied to align with the decoder’s output feature size. We obtain two losses, one is computed from the encoder features while the other one is computed from the decoder features. Consequently, the loss function is defined as follows:
$Loss=\alpha\times L_{enc}+(1-\alpha)\times L_{dec}$ (1)
where, $\alpha$ represents the contribution of the encoder loss in the objective function.
The DeCon-SL architecture is illustrated in Fig. 1 A, where $f_{\theta}$ and $f_{\epsilon}$ represent student and teacher encoders, respectively. $aux^{enc}_{\theta}$ and $aux^{enc}_{\epsilon}$ are the SSL framework’s auxiliary layers of the student and teacher encoders, respectively. Auxiliary layers can include any projector, predictor heads, or any layers used in the loss definition of the original SSL framework. $g_{\theta}$ and $g_{\epsilon}$ are the decoders of the student and teacher networks, and $aux^{dec}_{\theta}$ and $aux^{dec}_{\epsilon}$ their respective auxiliary layers. $g_{\theta}$ and $g_{\epsilon}$ take the output of encoder $f_{\theta}$ and $f_{\epsilon}$ as input, respectively. Auxiliary layers $aux^{enc}_{\theta}$ and $aux^{enc}_{\epsilon}$ are responsible for generating a representation from the encoder output, that is used to compute a contrastive loss for the encoders $L^{encoder}$. The contrastive loss, $L^{decoder}$ is calculated from the representation obtained from the auxiliary layers of the decoders, i.e., $aux^{dec}_{\theta}$ and $aux^{dec}_{\epsilon}$. $f_{\epsilon}$ and $g_{\epsilon}$ are updated using the exponential moving average (EMA) of $f_{\theta}$ and $g_{\theta}$, respectively.
3.2 DeCon-ML
To promote a comprehensive usage of the encoder’s parameters, we introduce a multi-layer contrastive decoder loss. This updated framework consists of a single loss for the encoder and multiple loss components for the decoder in a deep supervision manner along with channel dropout, which we describe in detail below. Figure 1 B illustrates the DeCon-ML adaptation of a given SSL framework.
Channel dropout: The dropout layer is typically used to prevent overfitting, improve generalization, and act as an implicit ensemble [50, 51]. However, we extend its use by applying dropout to channels transferred through skip connections between the encoder and decoder. In DeCon-ML, we apply channel dropout to the output layers of the different encoder levels by zero-ing out entire channels of the feature maps when passing these outputs to the decoder. No channel is zeroed-out when the inputs pass through the encoder. This channel dropout prevents the model from over-relying on specific features shared through skip connections. It encourages a more comprehensive use of the encoder’s parameters at the different levels, leading to a richer and more powerful data representation learned, while still preserving the input information content as it passes through the encoder to its bottleneck.
Decoder deep supervision: An encoder-decoder SSL framework allows us to retrieve meaningful features from multiple levels of the decoder. We can then pre-train the architecture with multiple losses using decoder deep supervision. If the decoder is connected to the encoder at multiple levels through lateral/skip connections, decoder deep supervision strengthens the representation power of the encoder at different levels [46]. Since a decoder passes the encoder bottleneck representation in a bottom-up manner to its upper level layers, the final encoder bottleneck representation remains a non-negligible objective of the training since it impacts all decoder level losses. In contrast, applying deep supervision at multiple encoder stages in an encoder-only SSL framework might prevent the creation of a strong semantic representation at the encoder bottleneck. To perform the decoder deep supervision, we add auxiliary layers to each decoder level, and a loss is computed at each level between the teacher and student representations. When performing decoder deep supervision, the encoder loss will remain the same as Eq. 1, but the decoder loss will be updated as follows:
$Loss=$ $\alpha\times L_{enc}+(1-\alpha)\times L_{dds}$ (2)
$L_{dds}=$ $\frac{1}{j}\times\sum_{i=1}^{j}L_{dec_{i}}$ (3)
where $j$ is the number of decoder levels at which a loss will be calculated and $L_{dec_{i}}$ is the loss computed at level $i$. The final decoder loss in Eq. 2 is thus an average of the losses at the different levels.
3.3 Implementation details
Architecture: We use a ResNet-50 encoder [26] as a backbone and implement a Fully Convolutional Network (FCN) [37] and a skip-connected Feature Pyramid Network (FPN) [34] as two decoder architectures in the DeCon framework. We follow the mmseg [14] FCN implementation from SlotCon [55] and customize the Detectron2 [56] FPN implementation. Details on these decoders are provided in Supplementary Material Section 1. We use DeCon to adapt the SlotCon [55] framework. Additional experiments are conducted using DenseCL [53] and PixPro [60] framework, or a ConvNeXt-Small [36] backbone.
For SlotCon, the student and teacher encoders, and every auxiliary layers—i.e., the respective projectors and semantic groupings, the slot predictor and the prototypes—are kept unchanged. Two decoders, a student and a teacher decoder (either FCN or FPN), that take the output of their corresponding encoder as input are added to the framework. The auxiliary layers are replicated at the decoder level. In the case of decoder deep supervision with FPN decoders (DeCon-ML), auxiliary layers are replicated independently at the four decoder levels. No deep supervision is performed in the DeCon adaptation of SlotCon with an FCN decoder (DeCon-SL). The decoder projectors differ from the encoder projectors only in input channels: 2048 for the encoder and 256 for the FPN and FCN decoders. We follow SlotCon by updating the student branch and the corresponding encoder and decoder slot predictors with backpropagation, the teacher branch with an EMA of the first one, and the prototypes centers with an EMA of the previous batch centers seen during training. Our implementation is based on the official SlotCon repository.
We adapt in a similar way DenseCL and PixPro into a DeCon-SL framework using an FCN decoder. More details are given in Supplementary Material Section 2. We adapt DenseCL implementation from mmselfsup package [15] version 1.0.0 and PixPro from the original implementation.
Although the pre-trained backbone remains unchanged, adding multiple decoder levels and associated loss functions increases the overall number of parameters in the DeCon-ML pre-training. To enable a fair comparison with the original SlotCon framework using a similar parameter budget, we implemented a reduced version of DeCon-ML that includes only the first two decoder levels originating from the encoder bottleneck. Additionally, the hidden dimension in the decoder projector was reduced from 4096 to 2048. This smaller version, which matches the parameter count of the original SlotCon framework (when $\alpha=0$), is referred to as DeCon-ML-S, while the full model with four decoder levels and unmodified configuration is denoted DeCon-ML-L. Detailed figures of DeCon-SL, DeCon-ML-S and DeCon-ML-L adaptations of SlotCon are available in the Supplementary Material Section 3.
Pre-training setup: We pre-train DeCon on COCO 2017 dataset [35], COCO+ (train2017 + unlabeled2017 datasets) and ImageNet-1K datasets [18]. The augmentations and hyperparameters—learning rate (LR), weight decay, schedulers, input shape, momentum, and so on—are kept the same as the adapted SSL framework. Models are pre-trained on COCO and COCO+ for 800 epochs, and for 200 epochs on ImageNet-1K. The SlotCon ResNet-50 models are pre-trained using a LARS optimizer, with a total batch size of 512 split across all available GPUs. The base LR is 1.0 scaled with the batch size following [55] and updated with a cosine LR decay schedule using a weight decay of $10^{-5}$ and 5 warm-up epochs. The ConvNeXt-Small backbone is pre-trained on COCO dataset with SlotCon and DeCon-ML using an AdamW optimizer [38], a base LR of 0.002 and weight-decay 0.05. To compare with ViT-based methods, we also pre-train a ConvNeXt-S backbone on ImageNet-1K for 250 epochs with DeCon-SL framework and an FCN decoder using a base LR of 0.001 and drop path rate of 0.1. Remaining hyperparameters are kept the same as for the ResNet-50 implementation. The pre-trainings using ConvNeXt-S backbones or FPN decoders are done using eight 80GB NVIDIA H100 GPUs. The pre-trainings without decoder and with FCN decoders are conducted on one 80GB NVIDIA A100 GPU.
To evaluate the generalization of our approach to different SSL frameworks, we adapt DenseCL and PixPro frameworks with DeCon-SL by pre-training a ResNet-50 encoder with an FCN decoder on COCO 2017 dataset. Optimization follows the base frameworks. To evaluate the dependence to the pre-training dataset, we also pre-train a similar DeCon-SL adaptation of SlotCon on REFUGE [41] and ISIC [13] datasets. We provide all pre-training details in Supplementary Material Section 2.
3.4 Evaluation protocol
We evaluate the different pre-trained models on various downstream tasks. Unless stated otherwise, we only transfer the pre-trained encoder for the fine-tuning experiments. We initialize the network with a pre-trained encoder and fine-tune end-to-end in all cases. All fine-tuning results reported are averaged over three independent runs. The pre-trainings, however, are only run once. All fine-tuning experiments related to the same downstream task are run on the same hardware. We perform semantic segmentation tasks using mmsegmentation [14] version 0.30.0. and other remaining tasks using Detectron2 package [56] version 0.6.
Object detection and instance segmentation: We fine-tune a Mask R-CNN with a pre-trained ResNet-50 or ConvNeXt-Small backbone on COCO 2017 dataset [35].
Semantic segmentation: We fine-tune a pre-trained ResNet-50 or ConvNeXt-Small with an FCN decoder on Pascal VOC [21], and ResNet-50 with an FCN decoder on ADE20K [63] and Cityscapes [16] following SlotCon [55] and PixCon [42]. We also fine-tune an ImageNet-1K-pre-trained ConvNeXt-Small backbone with an UPerNet decoder on ADE20K dataset.
Panoptic segmentation, keypoint detection, and dense pose estimation: We adapt task-specific Detectron2 configurations based on the SlotCon object detection architecture, and fine-tune on the COCO 2017 dataset for panoptic segmentation and keypoint detection, and on the COCO 2014 dataset for dense pose estimation.
Generalization to out-of-domain datasets: We fine-tune a ResNet-50-FCN architecture on REFUGE [41] and ISIC 2017 [13] datasets for medical semantic segmentation using 5, 25, and 100% of the training data. We also fine-tune a ResNet-50-deeplabv3+ architecture for PlantSeg  [54] semantic segmentation and a Faster R-CNN ResNet-50-FPN architecture for object detection on PlantDoc [49] and Detecting Disease  [47] datasets. Fine-tuning on agriculture datasets is performed using 10 and 100% of the datasets.
More details on the training hyperparameters and target datasets are provided in Supplementary Material Section 2.
4 Results
4.1 Experiments
Table 1: Performance of DeCon-SL ($\alpha=0.25$) and DeCon-ML (Small/Large, $\alpha=0$, dropout=0.5) adaptations of SlotCon with a ResNet-50 backbone pre-trained for 800 epochs on COCO/COCO+ and 200 epochs on ImageNet-1K, then fine-tuned on different downstream tasks. Only pre-trained encoders were transferred. PixCon ImageNet-1K results are not reported due to unavailable checkpoints. Results are averaged over three runs. Re-implemented, * Results from original paper, $\lozenge$ Result from SlotCon [55] paper.
Pret. Dataset Framework Backbone Pret. Dec. Object Det. Instance Seg. Semantic Seg. Dense pose Est. Panoptic Seg. Keypoint Det.
COCO COCO VOC City ADE COCO COCO COCO
AP AP50 AP75 AP AP50 AP75 mIoU mIoU mIoU GPS AP GPSm AP PQ AP
[…]
COCO PixCon (2024) [42]* ResNet-50 - 40.81 60.97 44.80 36.80 57.93 39.62 72.95 76.62 38.00 - - - -
SoCo-D (2024) [44]* - 40.3 60.1 44.0 35.1 56.9 37.6 - - - - - - -
SlotCon (2022) [55] - 40.81 60.95 44.37 36.80 57.98 39.54 71.50 75.95 38.57 63.39 64.84 40.31 65.66
DeCon-SL FCN 40.97 61.22 44.81 36.92 58.12 39.78 73.01 76.21 38.81 63.50 64.98 40.44 65.77
DeCon-ML-S FPN 40.97 61.20 44.71 36.94 58.20 39.63 72.80 76.21 38.36 63.75 65.08 40.52 65.66
DeCon-ML-L FPN 41.18 61.38 44.91 37.12 58.35 39.94 72.92 76.45 38.70 63.72 64.94 40.90 65.88
COCO+ PixCon (2024) [42]* ResNet-50 - 41.2 - - 37.1 - - 73.9 77.00 38.80 - - - -
SlotCon (2022) [55] - 41.63 62.10 45.67 37.57 59.07 40.45 73.93 76.43 39.11 - - - -
DeCon-SL FCN 41.86 62.43 45.73 37.75 59.40 40.48 74.46 76.65 39.25 - - - -
DeCon-ML-L FPN 42.08 62.42 46.13 37.84 59.41 40.75 75.36 77.00 39.04 - - - -
ImageNet-1K Supervised $\lozenge$ ResNet-50 - 39.7 59.5 43.3 35.9 56.6 38.6 74.4 74.6 37.9 - - - -
DINO (2021) [6] - 40.24 60.25 44.13 36.47 57.49 39.20 73.09 75.57 37.30 - - - -
SlotCon (2022) [55] - 41.69 62.07 45.59 37.59 58.97 40.49 75.02 76.15 38.97 - - - -
DeCon-ML-L FPN 41.80 62.12 45.73 37.73 59.08 40.68 75.40 76.51 39.01 - - - -
Table 2: Performance of DeCon-ML-L ($\alpha=0$, dropout = 0.5) and SlotCon with ConvNeXt-S encoder, pre-trained for 800 epochs on COCO and fine-tuned on segmentation and detection tasks.
Pret. Dataset Framework Pret. Dec. Obj. Det. Inst. Seg. Sem. Seg.
COCO COCO VOC
AP AP mIoU
COCO SlotCon - 44.07 39.67 73.24
DeCon-ML-L FPN 44.71 40.37 73.81
Table 3: Comparison of ViT and ConvNeXt models pre-trained on ImageNet-1K. We pre-trained a ConvNeXt-S encoder using DeCon-SL (FCN decoder, $\alpha=0.25$). All models are fine-tuned with an UPerNet decoder [58] on ADE20K. *Results from MixedAE [7]; †Results using two decoder levels to match with DeCon-SL decoder.
Framework Pret. Backb. Backb. Params. (M) Pret. Epoch ADE20k (mIoU)
MoCov3 (2021) [12]* ViT-B 88 600 46.8
DINO (2021) [6]* ViT-B 88 1600 46.9
BEiT (2021) [2]* ViT-B 88 300 44.7
MAE (2022) [24]* ViT-B 88 300 46.7
MixedAE (2023) [7]*† ViT-B 88 300 47.4
DeCon-SL [ours] ConvNeXt-S 50 250 48.02
ResNet-50 backbone: Table 1 and Supplementary Material Table S1 present the results of fine-tuning a ResNet-50 backbone pre-trained with various contrastive SSL frameworks. We mainly compare with SlotCon and PixCon as they demonstrated the most competitive performance in dense downstream tasks among previously published contrastive methods. DeCon-ML-L is consistently establishing a new SOTA for all tasks when pre-trained on ImageNet-1K and is consistently improving upon its base framework SlotCon and showing SOTA results for most of the evaluated downstream tasks when pre-trained on COCO and COCO+ dataset. Notably, it reduces the gap between PixCon and SlotCon for Pascal VOC and Cityscapes semantic segmentation when pre-trained on COCO. This supports the fact that the DeCon adaptation is robust across pre-training datasets and fine-tuning tasks. Moreover, DeCon-ML-S and DeCon-SL, whose training costs are comparable to SlotCon exhibit better performance than SlotCon and other contrastive frameworks. Finally, Tab. 1 also shows that performance improvement compared to SlotCon are robust to other downstream tasks such as dense pose estimation, panoptic segmentation and keypoint detection.
ConvNeXt backbone and comparisons with ViT: Table 2 shows the fine-tuning performance of a ConvNeXt-Small backbone pre-trained using DeCon-ML-L and the base SlotCon framework. A consistent improvement is seen with our encoder-decoder framework, confirming the robustness of our method to more modern and bigger backbones. Notably, both the absolute performance and the relative improvement are greater than those observed with the ResNet-50 backbone. Table 3 presents a performance comparison between a ConvNeXt-S model pre-trained on ImageNet-1K with DeCon-SL and ViT backbones pre-trained using various frameworks. Despite its smaller size and shorter pre-training schedule, the ConvNeXt-S backbone, pre-trained for dense tasks, outperforms all competing methods.
Table 4: Performance of DeCon-SL ($\alpha=0.5$) adapting DenseCL and PixPro with a ResNet-50 encoder, pre-trained for 800 epochs on COCO and fine-tuned on Pascal VOC and Cityscapes.
Pret. Dataset Framework Pret. Dec. Transfer Sem. Seg.
VOC City
mIoU mIoU
COCO DenseCL - Enc 71.11 75.87
DeCon-SL (DenseCL) FCN Enc 71.31 76.04
Enc + Dec 71.59 76.17
PixPro - Enc 72.13 75.84
DeCon-SL-(PixPro) FCN Enc 72.18 76.16
Enc + Dec 73.08 76.19
Generalizability across frameworks: Table 4 presents semantic segmentation fine-tuning results for models pre-trained with DenseCL and PixPro, along with their DeCon-SL adaptations. Pre-training with a decoder improves performance over the base frameworks, and transferring the decoder further enhances results, supporting the hypothesis that DeCon can be adapted to multiple frameworks. However, transferring the pre-trained decoder does not consistently benefit downstream performance in the case of SlotCon, as seen in Tab. 5 and Supplementary Material 5.
Out-of-domain tasks: Table 5 shows fine-tuning results of COCO-pre-trained encoders on REFUGE and ISIC. DeCon-SL pre-training consistently outperforms SlotCon, with larger gains under limited data. Transferring both encoder and decoder further improves performance on ISIC. Additionally, pre-training only on ISIC dataset consistently resulted in better performance compared to pre-training on COCO dataset. However, pre-training on REFUGE dataset did not outperform COCO pre-trained approaches, potentially due to the lack of available images (only 400) to develop a strong enough representation during pre-training. In both ISIC and REFUGE pre-trainings, DeCon-SL outperformed SlotCon. Table 6 shows that DeCon variants also perform better than SlotCon in agricultural object detection and semantic segmentation tasks using 10 or 100% of the datasets. DeCon pre-trained models also outperformed randomly initialized ones.
Table 5: Transfer performance with SlotCon and DeCon-SL (ResNet-50, FCN, $\alpha{=}0.5$) pre-trained on COCO, REFUGE, and ISIC, then fine-tuned on out-of-domain medical datasets (REFUGE and ISIC sem. seg. ) in various data settings.
SSL strategy Loss Transfer Pre-trained on COCO Pre-trained on Downstream Data
REFUGE ISIC REFUGE ISIC
mIoU mIoU mIoU mIoU mIoU mIoU mIoU mIoU
$L_{enc}$ $L_{dec}$ enc dec 5% 25% 100% 5% 25% 100% 100% 100%
Random init. - - - - 49.53 41.41 62.82 75.66 78.38 80.65 62.82 80.65
SlotCon 69.60 77.80 82.83 74.95 79.05 82.52 79.86 83.19
DeCon-SL 71.75 78.92 83.57 75.62 79.56 82.84 80.71 83.66
72.09 77.27 83.25 76.00 79.97 82.94 81.01 83.25
Table 6: ResNet-50 backbone transfer performance of SlotCon, DeCon-SL (FCN decoder) and DeCon-ML (FPN decoder), pre-trained on COCO and fine-tuned on agriculture datasets with 10% and 100% labeled data.
Framework PlantDoc (Obj. Det.) Detecting Diseases (Obj. Det.) PlantSeg (Sem. Seg.)
AP AP AP AP mIoU mIoU
10% 100% 10% 100% 10% 100%
Random Init. 6.59 19.81 19.14 34.76 16.96 24.96
SlotCon 17.53 38.37 26.54 48.45 20.72 28.78
DeCon-SL ($\alpha=0.25$) 17.06 39.94 26.81 48.85 20.56 29.39
DeCon-ML-S ($\alpha=0$, drop=0.5) 17.82 39.40 26.82 48.53 21.00 29.16
DeCon-ML-L ($\alpha=0$, drop=0.5) 17.98 40.03 27.16 48.57 20.66 29.77
4.2 Ablation studies
Table 7 presents an ablation study of the components in our DeCon-ML-L framework. The results show that while pre-training with a decoder and deep-supervision provides a modest benefit to downstream fine-tuning performance on COCO, the main performance improvement arises from the use of channel dropout in conjunction with decoder deep supervision. This combination yields the most substantial gains, highlighting the critical role of channel dropout when paired with deep supervision in our approach.
Table 7: Ablation of the different components of DeCon-ML-L and their impact on the downstream performance. $\Delta$ is the performance gap between the previous framework and the framework updated with the corresponding new component.
Framework Improvement COCO Obj Det. COCO Inst. Seg.
AP $\Delta$ AP $\Delta$
SlotCon 40.81 - 36.80 -
+FPN Decoder (Multi Level Loss) 40.85 +0.04 36.83 +0.03
+Dropout = 0.5 41.10 +0.25 37.05 +0.22
+Enc-Dec Loss Weight ($\alpha$) Tuning (DeCon-ML-L) 41.18 +0.08 37.12 +0.07
Table 8 shows DeCon-ML performance with different values of encoder loss weight $\alpha$, channel dropout, and number of decoder levels. A channel dropout probability of 0.5 provided the best downstream performance for DeCon-ML-L. Hence, this value was chosen to ablate $\alpha$. An $\alpha$ value of 0 gave the best downstream performance meaning that the decoder loss can replace the encoder one and pre-train efficiently an encoder on its own. We also ablated with number of decoder levels in DeCon-ML-L where 4 decoder levels yielded the best averaged performance across tasks. For DeCon-SL, an $\alpha$ value of 0.25 resulted in the best downstream performance, suggesting that having an encoder loss at the bottleneck is still beneficial to the pre-training in the absence of skip connections.
Table 8: Ablation of channel dropout, encoder loss weight ($\alpha$), and decoder levels in DeCon, pre-trained on COCO for 800 epochs and evaluated by fine-tuning on downstream tasks.
Abla tion Frame work Deco der $\alpha$ Drop. Obj. Det. Inst. Seg. Abla tion Frame work Deco der $\alpha$ Dec Levels Sem. Seg.
COCO COCO VOC City
AP AP mIoU mIoU
Dropout DeCon-ML-L (4 decoder levels) FPN 0.50 0 40.85 36.83 Dec. Level DeCon-ML-L (Drop = 0.5) FPN 0 2 72.70 76.07
0.25 40.95 36.97 3 73.14 76.10
0.50 41.10 37.05 4 72.92 76.45
0.75 41.02 37.05 - - - -
$\alpha$ 0.75 0.50 40.98 36.90 $\alpha$ DeCon-SL (No Drop) FCN 0.75 - 71.98 76.18
0.50 41.10 37.05 0.50 - 72.42 75.62
0.25 41.05 37.13 0.25 - 73.01 76.21
0 41.18 37.11 0 - 72.17 75.71
5 Discussion
In this paper, we introduced a contrastive framework adaptation by incorporating a decoder contrastive loss to an existing framework. We showed that our proposed SSL adaptation achieved SOTA performance in object detection, instance and semantic segmentation tasks when pre-trained on COCO+ and ImageNet-1K datasets, and in most tasks when pre-trained on COCO (Tab. 1). This adaptation enhances the representation of the pre-trained encoder, improving its transfer performance for dense downstream tasks, even in tasks beyond object detection, instance and semantic segmentation, regardless of the homogeneity between pre-training and fine-tuning decoders.
Figure 2: DeCon-ML-L and SlotCon pre-training loss dynamics.
We introduced a combined loss where encoder and decoder losses contribute to the total loss in a weighted manner. Figure 2 illustrates losses behavior across different pre-trainings when both the encoder and decoder are used. As shown, decoder pre-training does not affect the encoder loss dynamics, supporting the fact that the encoder and decoder loss terms are conceptually non-competing. We also observed that increasing the weight for the decoder loss compared to that of the encoder enhances the performance of the pre-trained models. This could be attributed to the non-competing nature of the encoder and decoder losses compared to approaches that use a reconstruction loss for a decoder and a different loss for the encoder in an encoder-decoder architecture [31]. Additionally, Fig. 3 presents the slots learned by SlotCon encoder and by the different decoder levels of DeCon-ML. It demonstrates that the decoder learns features that capture similar concepts as the encoder’s, while being more spatially precise.
Figure 3: Slots, as defined in SlotCon [55], learned at different outputs of our architecture. The top left image from COCO validation dataset is overlayed with the slot that was the most represented in original SlotCon’s encoder output feature map, resized to the image’s shape. Other slots displayed are the slots that had the biggest overlap with SlotCon’s main encoder output slot. More details on the slots creation are available in Supplementary Material 6.
Moreover, we noticed that the encoder loss is optional when pre-training with DeCon-ML, demonstrating the efficiency of the decoder deep supervision and channel dropout combination to ensure a comprehensive pre-training of both encoder and decoder. The lack of skip-connections in DeCon-SL with FCN decoder makes the combined encoder-decoder loss preferable to ensure the best pre-training. In Supplementary Material Section 7, we provide experiments on using DeCon-SL with an FPN decoder. We observed that with an FPN decoder, dropout does not provide the same benefits in DeCon-SL compared to DeCon-ML. Hence, channel dropout should only be applied to the different encoder levels when pre-training with a skip-connected decoder and decoder deep supervision (DeCon-ML). Framework-specific hyperparameter tuning of the decoder auxiliary layers could also increase the fine-tuning performance (Supplementary Material Section 8).
We demonstrated that our method is effective with different encoders. Specifically, pre-training a ConvNeXt-Small backbone, which is both more efficient and larger than the ResNet-50 backbone, also resulted in superior downstream performance when integrated with a decoder in the DeCon-ML-L framework. These results attest to the scalability and practical utility of the proposed approach for pre-training modern backbone architectures. Additionally, Tab. 3 shows that a ConvNeXt-S backbone pre-trained with DeCon-SL for dense tasks outperforms larger ViT-based SSL methods, despite using fewer pre-training epochs.
DeCon was also used to successfully adapt three different SSL frameworks using different concepts (pixel level VS object/group level) to learn local information (Tab. 4). This suggests that the proposed DeCon adaptation could be used in other SSL contrastive methods targeting dense tasks.
Table 9: Pre-training parameter count and computational cost for the different architectures. We study various pre-training setups on the COCO dataset using the SlotCon framework. Module sizes are given in millions of parameters. GPU-days are calculated on H100 80GB GPUs.
Pre-training Enc. size Dec. size Enc. Aux. Layers Size Dec. Aux. Layers Size Full Archi. Size GPU- days
SlotCon 23.51 x 2 0 71.51 0 118.52 1.53
DeCon-SL-FCN ($\alpha$ = 0.25) 23.51 x 2 5.31 x 2 71.51 56.83 185.97 1.77
DeCon-ML-S ($\alpha$ = 0) 23.51 x 2 3.35 x 2 0 59.08 110.03 1.78
DeCon-ML-L ($\alpha$ = 0) 23.51 x 2 3.35 x 2 0 227.31 281.02 11.00
DeCon adaptation of SSL frameworks yields clear advantages on small-scale datasets and out-of-domain dense prediction tasks, outperforming encoder-only baselines (Tab. 5, Tab. 6). These gains persist even when pre-training on the small target datasets, or when fine-tuning with only a fraction of labeled data. This highlights DeCon’s value in domains with limited annotations, like medical and agricultural applications.
Performance of deep learning models can vary across multiple runs due to inherent stochasticity. To mitigate the impact of randomness while considering computational resource constraints, each experiment was conducted three times, with all fine-tunings for a specific downstream task performed on similar hardware. The performance improvements achieved with our DeCon framework compared to the original framework are statistically significant, as evidenced by large Cohen’s $d$ values (>0.8>0.8) (in Supplementary Material Tab. S6), exceeding two standard deviations (see Supplementary Material Section 9), and $p$-values less than 0.05 (0.012 for COCO object detection and 0.048 for instance segmentation) from Wilcoxon signed-rank tests. Details on the computation of these values are provided in Supplementary Material Section 10. Furthermore, the performance gains, shown in Supplementary Material Tab. S6, are comparable to those reported in prior studies.
A limitation of our approach is its high memory and computational demand. However, as shown in Tab. 1, when GPU access is limited, reducing the number of decoder levels and tuning the decoder projector layers can decrease the parameter count and training cost of DeCon-ML (e.g., DeCon-ML-S in Tab. 9) while still enhancing pre-training performance. It is also notable that DeCon-SL and DeCon-ML-S have comparable training costs as their baseline.
Brempong et al. [5] followed a staged approach, first pre-training the encoder with supervision on annotated datasets, then pre-training the decoder with denoising. However, this multi-stage process relies on labeled data. It could also result in suboptimal feature alignment and increased training complexity. Future works could involve extending our proposed DeCon adaptation to offer a multi-stage continual pre-training in a contrastive manner. We also plan to adapt DeCon to ViT architectures in the future.
6 Conclusion
In this paper, we introduced DeCon—a novel contrastive framework adaptation that enhances the performance of SSL approaches for various dense predictions tasks. We showed that this adaptation improves the encoder’s representation significantly. We proposed two variants of DeCon: DeCon-SL and DeCon-ML. The former introduces a decoder contrastive loss, and the latter extends DeCon-SL by introducing channel dropout and decoder deep supervision to maximize the encoder’s pre-training power. By integrating the proposed contrastive joint encoder-decoder pre-training strategy into an existing framework, we show that DeCon consistently outperforms or matches state-of-the-art methods across diverse dense prediction tasks, highlighting the effectiveness of this unified pre-training approach.
Acknowledgments
We extend our sincere gratitude to the University of Calgary VPR Office for their generous support through VPR Catalyst Grants, which was instrumental in making this research possible. We also acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC Discovery Grant RGPIN-2024-04966) and Alberta Innovates. This work was partly supported by Alberta Innovates Graduate Student Scholarship and T. Chen Fong Doctoral Research Excellence Scholarship in Imaging Science. Computing resources were provided by the University of British Columbia Cloud Innovation Centre, powered by Amazon Web Services https://cic.ubc.ca/. This research was also partly enabled by support provided by the Digital Research Alliance of Canada https://alliancecan.ca/, and the Research Computing Services group at the University of Calgary https://rcs.ucalgary.ca/RCS_Home_Page.